### 25 Mar 2026 v2

# VHE gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022

Abstract
Context. BL Lacertae (BL Lac), the archetype blazar of its subclass and one of the most studied blazars in the last decades, has gone through a series of major multi-wavelength outbursts since 2020, resulting in its highest recorded -ray flare up to date between September and November 2022 together with those from August 2021 and October 2024.
Aims. We characterise the -ray and multi-wavelength emission and spectral energy distribution (SED) of BL Lac, as well as their evolution during the major and extended -ray and multi-wavelength flare occurring between September and November 2022.
Methods. We evaluate the variability of the flare, with focus on the nights of October 20 and November 13, when clear intranight very-high-energy (VHE, GeV) -ray variability is observed. We model the -ray and broadband SEDs during periods of stable emission identified with a Bayesian block analysis, interpreting their evolution of the flare from the variability of the relativistic particles and physical parameters of the jet.
Results. The VHE emission shows an average flux of 0.23 Crab Units (C.U.) above 200 GeV during this flare and a variability amplitude of more than a factor 10. Intranight doubling-flux variations as fast as 8 minutes are observed during the nights of October 20 and November 13, when maximum fluxes of 4.4 C.U. above 100 GeV and 2.8 C.U. above 200 GeV are reached. The spectral analysis reveals a transition of the X-ray emission from the high- to the low-energy SED peak, and a shift of the -ray peak towards higher energies. The broadband emission was interpreted within a leptonic two-zone model in which intranight variability is explained as magnetic reconnection in a compact region closely oriented with the line of sight, while weekly-scale variations can be explained as variations of the relativistic electron distributions and the injection of freshly accelerated particles.
Key Words.:
Galaxies: active – BL Lacertae objects: general – BL Lacertae objects: individual: BL Lacertae – Galaxies: jets – Galaxies: nuclei – gamma rays: galaxies1 Introduction
Relativistic jets launched from blazars — active galactic nuclei (AGN) with their jets oriented towards our line of sight — are known to be extremely efficient and natural particle accelerators, which are able to emit in all the electromagnetic spectrum. In fact, they are one of the few sources that can emit up to TeV energies. Blazars are also highly variable on all timescales. Depending on the properties of their optical spectrum (absence/presence of broad emission lines, see Stickel et al., 1991), they can be classified into two main sub-classes: BL Lac type objects and flat spectrum radio quasars (FSRQs) (Urry & Padovani, 1995).
The broadband spectral energy distribution (SED) of blazars, extending from radio to -ray energies, has a characteristic double-peak shape. The low-energy component is produced by synchrotron radiation (see Konigl, 1981). The invoked mechanisms behind the high-energy component fall in two groups. An interpretation based on leptonic processes explains this emission through Inverse Compton (IC) scattering with low-energy photons, either synchrotron radiation via synchrotron self Compton (SSC, see Konigl, 1981; Maraschi et al., 1992) scattering, or with radiation from outside the jet via External Compton (EC, see Dermer & Schlickeiser, 1994) scattering. However, interpretations based on a hadronic nature are also used in some cases, relying on processes like proton synchrotron radiation or photo-pion production to explain the high-energy emission of blazars (see e.g. Mannheim & Biermann, 1992; Aharonian, 2000).
BL Lacertae (hereafter BL Lac), the archetype of the BL Lac blazar subclass located at redshift (Miller et al., 1978), has historically shown remarkable variability with several episodes of minute- to hour-scale variations (see for instance Abeysekara et al., 2018; MAGIC Collaboration et al., 2019; Jorstad et al., 2022), becoming one of the most monitored and studied blazars over the years. It is classified as an intermediate-synchrotron-peaked (ISP) blazar according to the peak frequency of its synchrotron emission (). However, it is known for transitioning between the low-synchrotron-peaked (LSP, ) and ISP types, depending on its activity (see e.g. Ackermann et al., 2011; Nilsson et al., 2018). It has been detected several times in the very-high-energy (VHE, GeV) -ray band, always during multi-wavelength flaring emission states (Albert et al., 2007; Arlen et al., 2013). These periods of variability on very short timescales imply very compact emitting regions involving the most energetic particle acceleration processes, and challenging in some cases the theoretical models used for interpreting the broadband emission of blazars (Henri & Saugé, 2006).
Here we report the results of an extreme flare observed from BL Lac by the Large-Sized Telescope prototype (LST-1), taking place between September and November 2022, and triggered after the detection of enhanced activity by the Fermi satellite (La Mura, 2022). This flare is one of the three brightest events ever detected in VHEs from BL Lac, together with similar events occurring in 2021 and 2024, and it is also the most long-lasting flaring period ever recorded for this source in the VHE domain, lasting for about 3 months, with significant variations in day, hour and minute timescales. This remarkable flare was also covered by several multi-wavelength facilities, providing a rich dataset for interpreting the multi-wavelength emission and variability of BL Lac in this period (see Sect. 2). In this work, we evaluate and interpret the VHE -ray and broadband evolution of the flare. We characterise the minute-scale -ray variability observed during this period (Sect. 3), that sets important constrains to the nature of the emitting region. The spectral behaviour of the source is evaluated (Sect. 4), and its broadband emission is modelled within AGN radiative models (Sect. 5). The main results are discussed and summarised in Sects. 6 and 7.
2 Observations and data reduction
2.1 VHE -ray data: LST-1
LST-1 is the first of the four Large-Sized Telescopes that will be constructed in the northern site of the Cherenkov Telescope Array Observatory (CTAO, Acharyya et al., 2019) at the Roque de los Muchachos Observatory in the Canary Island of La Palma (Spain). LST-1 is currently performing commissioning and scientific observations, after the completion of its construction in October 2018. Its large collection area (23 m diameter mirror) and very sensitive camera — composed of 1855 high quantum efficiency photomultiplier tubes — allow to detect faint Cherenkov light flashes from -ray showers down to energies 20 GeV (see Abe et al., 2023).
LST-1 observed BL Lac for 23 nights between September 21 and November 25, 2022 in wobble observing mode (Fomin et al., 1994), providing 36 hours of data after applying standard data quality cuts in a zenith range between 10∘ and 60∘. The analysis of the raw LST-1 data was performed with version v0.10.18 of the cta-lstchain software (López-Coto et al., 2022; Moralejo et al., 2025), including data cleaning, image parametrisation, -hadron separation and direction and energy reconstruction of the recorded events. Among this dataset, 29 hours of data were taken under dark sky conditions (no increased background light contamination due to the presence of the moon) with a standard tail-cut cleaning with thresholds of 8 and 4 photoelectrons (p.e.) for the core and neighbour pixels of the Cherenkov shower image and an intensity cut of 50 p.e. (Abe et al., 2023). The remaining 7 hours of observation were performed with different moon brightness levels. These data were analysed with cleaning thresholds ranging from 10 to 16 p.e. and 5 to 8 p.e. for the core and neighbour pixels, and increased intensity cuts between 63 and 194 p.e. due to the noise introduced by the moon, extracted from the data at each night sky background (NSB) level. The event classification was performed with a Random Forest trained using Monte Carlo (MC) simulated proton- and -like events at a declination of 34.74∘, close to that of BL Lac (42.28∘). The MCs were produced to match the NSB level of our data, accounting for the additional noise with respect to dark observations. For the event selection, energy-dependent gammaness (gamma-likeness score) and (angular distance of the detected event with respect to the nominal source position) cuts were used, optimised to preserve 70% of the simulated MC -ray events.
The high level analysis was done with version 1.2 of Gammapy (Donath et al., 2023; Aguasca-Cabot et al., 2023; Acero et al., 2024). The background was characterised using an OFF region with a wobble angle of 0.4∘ and at a reflected position of 180∘ from that of BL Lac. A limit on the effective area ¿5% of the maximum area was also applied. The statistical significance was calculated following Eq. (17) from Li & Ma (1983). The energy threshold of the light curve, calculated from the maximum of the distribution of expected events from the MCs, weighted with BL Lac’s spectrum, with the highest NSB level and zenith distance between 40∘ and 50∘ to match that of the data, was estimated to be 175 GeV. Therefore, we have used a minimum energy of 200 GeV for the light curve reconstruction. The energy threshold was further verified on data against background asymmetries in the field of view (FoV). We observed a non-uniformity of the event distribution, in particular along the altitude axis. This effect increases when the telescope pointing is further from the zenith, and becomes clearly visible at zenith distances ¿40∘. Looking at the energy distribution of events in the source region and the OFF region, a clear difference in energy threshold is visible for some observations. While the difference seems mostly at energies below our analysis threshold, we also perform our spectral analysis using backgrounds counts estimation using the package BAccMod111https://github.com/mdebony/BAccMod. This package allows to extract the background distribution using the full FoV after removing the source region and can thus account for asymmetries. As expected, we observe differences in the extracted spectra below 200 GeV for some time periods, but the spectra agree above 200 GeV thus validating our results above this energy. The VHE -ray light curve is shown in the top panel of Fig. 1. We highlight October 20 and November 13 as two remarkably bright nights, when significances of 49 and 30 were reached in only 2.5 and 1.5 hours of observation. The minute-scale variability of these two nights is evaluated in Sect. 3. In particular for October 20, consisting only of dark-sky data, we adopted a threshold of 100 GeV obtained from the dark-sky MCs with zenith 20-40∘. The spectrum and SED were characterised with a log-parabola model plus a term accounting for the Extragalactic Background Light (EBL) absorption with the model from Saldana-Lopez et al. (2021). The spectral analysis is presented in detail in Sect. 4.
2.2 HE -ray data: Fermi-LAT
We also analysed high-energy (HE, MeV) data taken by the pair-conversion Large Area Telescope (LAT) on board the Fermi satellite. The LAT operates in sky-survey mode in an energy range between 20 MeV and 1 TeV, with its best sensitivity around a few GeV (Atwood et al., 2009). We considered a 20∘-radius region of interest (ROI) centred at the position of BL Lac. All data from September 21 to November 30 were downloaded, selecting the Pass8 P8R3_SOURCE_V3 events between 100 MeV and 300 GeV with the Fermitools standard software. A zenith angle cut of 90∘ was applied to minimise the contamination from the limb of the Earth. Furthermore, we used the recommended versions of the Galactic diffuse emission model and isotropic component for the event selection considered222https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html.
To account for the emission of sources nearby BL Lac, we derived a model including all sources within the ROI, plus all sources located in an additional annular region of radius 10∘. This model was constructed using 1 year of data and a binned likelihood analysis procedure, using as reference the 4FGL-DR2 catalogue (Abdollahi et al., 2020; Ballet et al., 2020). The spectral parameters (index and normalisation) of all variable sources — as defined in the 4FGL catalogue based on the LAT variability index — contained in the ROI and with a Test Statistics (TS, Mattox et al., 1996) ¿25 (5 significance) were left as free parameters, while those not fulfilling these conditions were fixed to the catalogue values. The normalisation of the diffuse components was also left free. Convergence to the final model was achieved after removing all sources with TS¡4 (¡2).
The HE -ray light curve was then obtained analysing the data within the desired time window through an unbinned likelihood analysis of events contained in 12-hour bins. For each bin we have assumed a power law spectral shape for BL Lac, leaving the spectral index and the normalisation free. The spectral parameters of the rest of the sources were fixed to the values from the model, with the exception of two highly variable and bright sources with a significant detection (TS¿25) located at a distance ¡15∘ from BL Lac (4FGL J2244.2+4057 and 4FGL J2311.0+3425). The HE -ray light curve is shown in Fig. 1.
2.3 X-ray data: Swift-XRT
The scientific payload onboard the Neil Gehrels Swift observatory includes by the X-Ray Telescope (XRT, Burrows et al., 2005), and the Optical/UV Telescope (UVOT, Roming et al., 2005). In standard operation mode, these instruments perform simultaneous observations on the same targets. In this project, we included 50 Swift-XRT and Swift-UVOT observations on BL Lac carried out since March 2022 up to January 2023.
Swift-XRT is sensitive to X-rays in the energy range. The reduced XRT data were downloaded from the UK Swift Science Data Center (Evans et al., 2009). The spectral analysis on the reduced data was performed using the XSPEC v12.13.1 package through PyXspec v2.1.2. We analysed the XRT data in the energy range from 0.3 to 10 keV. The intrinsic source emission in this range was primarily modelled with a power-law function and, to test the presence of intrinsic spectral curvature related to the transition between the synchrotron and IC bumps, with a log-parabola model. The observed spectra were fitted with the intrinsic emission models folded with the tbabs333https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node273.html Galactic absorption term, as implemented in XSPEC. Following Raiteri et al. (2009), in the fit procedures we fixed the neutral hydrogen column density toward BL Lac to and we employed this value to correct the X-ray fluxes and SEDs for the Galactic extinction effects. The test was employed to assess whether the log-parabola model was preferred over the power-law one, with a significance level cut. The results of the X-ray analysis are summarised in Appendix B.
2.4 Ultraviolet data: Swift-UVOT
The analysis of UVOT data was carried out using the uvotimsum and uvotsource tasks within the HEAsoft 6.32 package, along with the latest (v20240201) Swift/UVOT Calibration Database (CALDB) files. UVOT performed photometric observations on BL Lac using the optical vv, bb, uu and UV w1, m2, w2 filters (Poole et al., 2008). Through uvotsource, we estimated the observed fluxes in the UVOT filters using a circular source region with radius of 5 centred on BL Lac and a circular background region with a 10 radius in a nearby region devoid of sources. The observed fluxes were corrected for Galactic extinction effects using (Schlafly & Finkbeiner, 2011) using the standard procedure described in Fitzpatrick (1999). Moreover, the contributions of the host galaxy in the UVOT photometric filters vv, bb, uu, w1, m2, w2, given the adopted aperture, were estimated as 2.89, 1.30, 0.36, 0.026, 0.020, and 0.017 mJy, respectively (Raiteri et al., 2013). Both host galaxy contributions and Galactic extinction effects were accounted for to reconstruct the intrinsic BL Lac spectral points in the optical-UV band. The results of the analysis of the Swift-UVOT data from the 2022 high state are reported in Table 7.
2.5 Optical data
Optical observations in the Johnson-Cousins bands were taken by the Sierra Nevada Observatory (SNO) as part of the TOP-MAPCAT photo-polarimetric blazar monitoring programme (Agudo et al., 2012). The data were reduced with the IOP4 software (Escudero Pedrosa et al., 2024a, b). Observations in the band were also performed as part of the Tuorla blazar monitoring444https://tuorlablazar.utu.fi/ with the 80 cm Joan Oró Telescope at Montsec Astronomical Observatory (Spain), reduced as detailed in Nilsson et al. (2018). Data in the bands were taken by the IAC80 telescope at the Teide Observatory. Finally, observations with the robotic telescope network of Las Cumbres Observatory (Brown et al., 2013) were also coordinated in the optical bands.
Additionally, data from the public all-sky survey programmes and databases All-Sky Automated Survey for Supernovae (ASAS-SN, Shappee et al., 2014; Kochanek et al., 2017) in the Sloan band, and Zwicky Transient Facility (ZTF, Bellm et al., 2019) in the Sloan bands, were also included. These Sloan magnitudes were converted into the Johnson-Cousins photometric system using the transformations from Lupton (2005)555https://classic.sdss.org/dr4/algorithms/sdssUBVRITransform.php#Lupton2005.
The contribution of the host galaxy emission to the total observed optical emission contained in the employed aperture has been subtracted (see Raiteri et al., 2009). Moreover, all data have been corrected from the reddening effect introduced by the Galactic extinction. For this, we adopt the extinction values , , and reported by Schlafly & Finkbeiner (2011). The optical flux density light curves are presented in Fig. 1.
2.6 Radio-millimetric data
Observations in the 37-GHz radio band were performed with the 13.7-m diameter Metsähovi radio telescope as part of its long-term AGN monitoring666https://www.aalto.fi/en/metsahovi-radio-observatory/active-galaxies (Teraesranta et al., 1998). Additionally, we compiled five observations in the 225-GHz band performed with the Submillimeter Array (SMA, Ho et al., 2004) at Mauna Kea, Hawaii. SMA is an 8-element 6-meter diameter dish radio interferometer with two orthogonally polarised receivers, allowing us to measure the total and the polarised flux of astrophysical objects. These receivers are inherently linearly polarised, however they can be converted to circular thanks to quarter-wave plates installed on the polarimeter (Marrone & Rao, 2008). These observations were carried out by the SMA Monitoring of AGNs with Polarisation (SMAPOL) programme (Myserlis et al., 2025), dedicated to the regular monitoring of -ray loud blazars, including BL Lac, on a bi-weekly cadence. The data were calibrated with the MIR777https://lweb.cfa.harvard.edu/~cqi/mircook.html package following standard procedures.
3 VHE intranight variability
In addition to the day-to-week variability, we have evaluated the VHE -ray intranight variability, a feature less often observed in blazars (other examples are e.g. Aharonian et al., 2007; MAGIC Collaboration et al., 2020). Its detection presents significant constrains to the emitting region size and the interpretation of the emission, potentially favouring leptonic models as hadrons usually require longer times for being accelerated and have longer cooling times by a factor . We evaluate the nights with the highest VHE -ray flux, October 20 and November 13. The 5-minute binned light curves are shown in Fig. 2.
3.1 October 20th flare
We have evaluated the fastest variability timescales observed during the night of October 20. For this night we consider the integrated flux above an energy of 100 GeV, as all the data were taken under dark-sky conditions and zenith distances 20-40∘. As shown in the left panel of Fig. 2, the rise and decay of the flare is well sampled by the LST-1 data, varying from 1.0 C.U. to 4.4 C.U. in hours. This well defined rise and decay profile of the flare during this night enabled us to model the light curve as a rising and decaying exponential function, commonly used for modelling light curves (Norris et al., 1996) including those from blazar flares (see e.g. Aharonian et al., 2007),
where represents the flux of the flare at a time , is the baseline flux that we have fixed at cm-2 s-1, obtained from the stable bins at the beginning and end of the flare, and and are the timescales of the flare rise and decay variations by a factor of . We have modelled the 5-minute binned light curve of October 20 as a sum of three possible substructures during the development of the flare with peak values found at min, min and min, considering the first flux point as min. We note that we have also considered the simpler cases of one or two components, comparing the goodness of the fit in each case. We measure values for the triple-, double- and single-exponential profiles of , and , which yields a significance of the triple-exponential model over the double- and single-function of and , respectively, using Wilks theorem (Wilks, 1938). We also checked that the light curve modelling does not improve adding a fourth component, with a preference over the three-component one of 0.2. Therefore, we adopted the model based on the sum of three exponential profiles defined by Eq. (1).
The fitted function is shown in the left panel of Fig. 2, and the calculated rise and decay timescales of exponential profile are reported in Table 1. These timescales can be used to estimate the doubling flux variability timescale during the rise and decay, representative of the fastest significant variability timescales, as
With this approach we are able to resolve variability timescales of 18 min with a confidence level of ¿, and faster hints of between 2.2 min and 4.6 min, however with a significance 2.
| Date | ||||||||
| [10-9 cm-2 s-1] | [10-9 cm-2 s-1] | [min] | [min] | [min] | [min] | [min] | [min] | |
| October 20 | ||||||||
| November 13 | 0 | – | – | – |
Alternatively, we can also estimate the fastest variability timescales through the doubling time criterion considering flux variations between consecutive temporal bins, following the procedure from Zhang et al. (1999) as
where and are the fluxes at times and . We consider only those timescales obtained through this method for which the significance of the variability is 2. The results derived with this approach are reported in the last column of Table 1. We observe significant doubling flux timescales of min.
Applying causality arguments, it is expected that very fast variability originates in very compact jet regions. Therefore, the fastest variability timescales observed can be used to constrain the size of the emitting region responsible for such variations as
where is the radius of the emitting region, represents the fastest variability timescale observed, is the speed of light, is the Doppler factor of the region and corresponds to the redshift of the source (see Dondi & Ghisellini, 1995; Tavecchio et al., 2010). Here we adopt a conservative approach, using the shortest timescale with a significance ¿ before trials, this is, min. With this timescale, and considering a range of Doppler factor values , the size of the emitting region responsible for the fast variability must be cm to cm. We note that if the 2.2- to 6.7-minute timescales are confirmed, this would imply a region a factor 1.5 to 4 more compact. Nevertheless, in addition to the low significance, when considering the typical systematic errors for LST-1 flux measurements (see Abe et al., 2023), these timescales get hindered within uncertainties.
3.2 November 13th flare
The same analysis was performed for the night of November 13, when we also observed intranight variability in the VHE -ray emission of BL Lac. In this case, and owing to the higher energy threshold introduced by the moon conditions, we consider the flux above 200 GeV. During this night, the emission decayed from 2.8 C.U. to 1.0 C.U. in 1.5 hours. The corresponding 5-minute binned light curve is shown in the right panel of Fig. 2. In this case, LST-1 covered only the decay of the VHE -ray flare, that we modelled with a decaying exponential function as
that follows the same description as Eq. (1). A value of cm-2 s-1 was used for , where the VHE flux seems to stabilise after the flare. The results for this night are reported as well in Table 1. With the estimated timescale min, we have again constrained the size of the emitting region with Eq. (4), for typical Doppler factors . This leads to a size constraint ranging from cm to cm. We note that in this case, the doubling flux criterion does not find significant minute-scale variability.
4 Characterisation of the HE-VHE -ray spectrum
| Period | ||||||
| [MJD] | [GeV-1 cm-2 s-1] | [GeV] | [erg cm-2 s-1] | |||
| BB1 | 59843.5 – 59856.5 | 1.4 | ||||
| BB2 | 59867.5 – 59870.5 | 1.9 | ||||
| BB3 | 59870.5 – 59872.5 | 1.9 | ||||
| BB4 | 59872.5 – 59873.5 | 5.5 | ||||
| BB5 | 59873.5 – 59879.5 | 1.1 | ||||
| BB6 | 59879.5 – 59880.5 | 2.3 | ||||
| BB7 | 59881.5 – 59885.5 | 1.6 | ||||
| BB8 | 59896.5 – 59897.5 | 6.5 | ||||
| BB9 | 59898.5 – 59903.5 | 1.9 | ||||
| BB10 | 59908.5 – 59909.5 | 2.3 |
In order to evaluate possible -ray spectral variability patterns, we have studied the evolution of the -ray spectrum during the flare. For this, we have applied a Bayesian Block (BB) analysis on the VHE -ray light curve using the algorithm developed by Wagner et al. (2022)888https://github.com/swagner-astro/lightcurves, identifying 10 BBs of comparable -ray activity with a false positive rate of 1% (see Fig. 1). Then, making use of the LAT and LST-1 data, we characterise the high-energy SED peak for the different identified periods from 100 MeV up to a few TeV. For this, we perform a joint fit of the data from both instruments with Gammapy, that allows to combine multi-instrument data at an event level, correctly accounting for instrument response functions (IRFs), backgrounds and other effects such as EBL absorption (Nievas Rosillo et al., 2025).
For the Fermi-LAT data of each BB, we filter out all sources in the model that have a TS¡10. The parameters of all sources contained in a circular region of 10∘ radius centred at the position of BL Lac and with a TS¿30 are left free, while for those sources within 5∘ from BL Lac with TS¡30, only the spectral normalisation is left as a free parameter. For the remaining sources, all parameters are fixed to the catalogue values. Then, a pre-fit of the Fermi-LAT data is performed, allowing to remove weak sources still included in the model for which the number of expected counts is ¡5. Finally, we combine the LAT and LST-1 datasets of each BB and calculate the complete -ray spectrum and SED peak, incorporating into the analysis the response of each instrument.
We have tested several spectral shapes of the HE-VHE emission, including log-parabola, power-law with exponential cut-off and log-parabola with exponential cut-off spectral models. We find that the first function provides the most accurate fit to our data. Therefore, we use a model consisting on a log-parabola function plus a term accounting for the EBL absorption at redshift with the model from Saldana-Lopez et al. (2021),
To perform a direct comparison between the different spectra, we adopted an GeV for all periods. The best-fit spectral parameters are reported in Table 2. In Fig. 3 we show an example of the joint SED for the eighth BB, that is, the night of November 13. The rest of the SEDs are available in Figs. 6 and 7. Alternatively, owing to the very small variability in the spectral curvature, we tested a log-parabola function fixing , taken from the average value over the entire period. The results were consistent within uncertainties with those reported in Table 2.
The complete characterisation of the high-energy SED peak performed in Sect 4 allows us to search for spectral variability patterns during the evolution of the flare. We observe a relatively stable value of the spectral index at 1 GeV around , with almost all BBs consistent within errors at a 2 level. Nevertheless, spectral index variations are particularly noticeable for the nights of October 20 with , and November 13 with , meaning that during these flares, particles are being accelerated to much higher energies. This is also visible from the high-energy peak shift, reported in Table 2. The IC peak shift becomes evident for the four brightest BBs, this is, the nights of October 20 and 27, and November 13 and 25, being the only ones with GeV. The average value of is found below 2 GeV, and during these nights we observe shifts up to values of 5.5 GeV for October 20, 6.5 GeV for November 13, and 2.3 GeV for October 27 and November 25. The Spearman correlation coefficient between and , and and show a high level of correlation ( and , respectively), indicating a possible harder-when-brighter behaviour. However, this correlation coefficient is dominated by the two brightest nights and does not account for uncertainties.
5 Broadband SEDs
We have modelled the evolution of the broadband emission of BL Lac during the 2022 high state. We have considered 10 broadband SEDs, corresponding to each of the BBs identified above. Fig. 4 shows as an example the model for the eighth BB. All fitted SED models are presented in Figs. 8 and 9. We note that, due to the lack of strictly simultaneous Swift data (0.5 days with respect to the LST-1 observations) to BBs 3, 4, 6, 8 and 9, for these periods we have considered a window of 1.5 days in order to have complete multi-wavelength coverage of the SEDs.
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) | (12) | (13) | (14) |
| Epoch | Component | B | K | R | |||||||||
| () | () | () | [G] | [ cm-3] | [ cm] | [min] | |||||||
| BB1 | Blob | 3.00 | 0.90 | 0.17 | 2.0 | 2.8 | 0.50 | 3.00 | 0.3 | 33 | 33 | 35.0 | |
| Core | 0.27 | 0.20 | 0.05 | 2.2 | 3.4 | 0.20 | 0.01 | 100 | 20 | 20 | 1.64 | – | |
| BB2 | Blob | 2.50 | 1.00 | 1.50 | 2.8 | 3.8 | 0.50 | 3.00 | 0.3 | 33 | 33 | 15.3 | |
| Core | 0.30 | 0.23 | 0.05 | 1.9 | 2.9 | 0.20 | 0.01 | 100 | 20 | 20 | 1.54 | – | |
| BB3 | Blob | 3.00 | 0.95 | 0.80 | 2.9 | 3.7 | 0.50 | 3.00 | 0.3 | 33 | 33 | 19.2 | |
| Core | 0.25 | 0.23 | 0.05 | 2.0 | 2.9 | 0.20 | 0.01 | 100 | 20 | 20 | 0.75 | – | |
| BB4 | Blob | 2.50 | 0.95 | 1.80 | 2.6 | 3.5 | 0.50 | 5.00 | 0.3 | 33 | 33 | 11.1 | |
| Core | 0.30 | 0.19 | 0.07 | 2.0 | 2.8 | 0.20 | 0.01 | 100 | 20 | 20 | 2.00 | – | |
| BB5 | Blob | 3.20 | 1.00 | 0.35 | 2.8 | 3.8 | 0.50 | 3.00 | 0.3 | 33 | 33 | 17.9 | |
| Core | 0.25 | 0.20 | 0.07 | 2.4 | 3.4 | 0.20 | 0.01 | 100 | 20 | 20 | 0.79 | – | |
| BB6 | Blob | 2.00 | 0.95 | 1.10 | 2.7 | 3.4 | 0.50 | 3.00 | 0.3 | 33 | 33 | 16.2 | |
| Core | 0.50 | 0.25 | 0.07 | 2.4 | 3.7 | 0.20 | 0.01 | 100 | 20 | 20 | 0.88 | – | |
| BB7 | Blob | 2.70 | 1.05 | 1.00 | 2.7 | 3.6 | 0.50 | 3.00 | 0.3 | 33 | 33 | 12.3 | |
| Core | 0.30 | 0.23 | 0.07 | 2.2 | 3.3 | 0.20 | 0.01 | 100 | 20 | 20 | 1.19 | – | |
| BB8 | Blob | 4.00 | 1.00 | 3.80 | 2.1 | 3.8 | 0.50 | 7.50 | 0.3 | 33 | 33 | 8.8 | |
| Core | 0.30 | 0.18 | 0.07 | 1.9 | 3.0 | 0.20 | 0.01 | 100 | 20 | 20 | 2.56 | – | |
| BB9 | Blob | 3.60 | 1.00 | 0.80 | 2.7 | 3.6 | 0.50 | 3.00 | 0.3 | 33 | 33 | 17.5 | |
| Core | 0.15 | 0.19 | 0.06 | 2.3 | 2.7 | 0.20 | 0.01 | 100 | 20 | 20 | 0.85 | – | |
| BB10 | Blob | 2.00 | 1.05 | 1.00 | 2.6 | 3.6 | 0.50 | 3.00 | 0.3 | 33 | 33 | 12.1 | |
| Core | 0.28 | 0.20 | 0.06 | 2.3 | 3.0 | 0.20 | 0.01 | 100 | 20 | 20 | 1.45 | – |
We have considered a leptonic-only origin for the broadband emission. We use version v0.4.0 of the python package agnpy (Nigro et al., 2022, 2023). We find that SSC models based on a single emitting region are not able to accurately describe the broadband emission of BL Lac during this period. One-zone models based on EC scattering rise as an alternative to explain the VHE -ray emission, and have actually been invoked in the past for BL Lac due to the presence of a small and weak broad line region (BLR), using its external low-energy photons as seeds for the -ray production (MAGIC Collaboration et al., 2019). However, this solution was also found to be insufficient in this case to reproduce the broadband emission and variability during the 2022 high state, especially given the tight constrains on the emitting region size set by the intranight variability. Therefore, we consider a SSC model based on two emitting regions (Tavecchio et al., 2011), which has been used frequently to model the multi-wavelength emission of this source (see e.g. MAGIC Collaboration et al., 2019; Sahakyan & Giommi, 2022). We denote the two regions as ”core” and ”blob”, with sizes . We consider the regions to be co-spatial and interacting. This is motivated by the high correlation of the broadband emission from the optical band up to rays (Appendix D.2). A scheme of this layout is illustrated in Fig. 9b of MAGIC Collaboration et al. (2019). Each region is defined by a magnetic field , a Lorentz factor , a Dopper factor and an electron population that we have characterised through a broken power-law distribution with respect to the electron Lorentz factor ,
where is the normalisation of the distribution, and are the minimum and maximum values of the Lorentz factor, and and are the spectral indices below and above the break of the distribution , respectively. We constrain the size of the blob, responsible for the intranight VHE -ray variability, with the estimates presented in Sect. 3. Additionally, we have incorporated to the SED model an EC contribution from external photon fields, namely the BLR, the accretion disc and the dusty torus. This is because, as mentioned above, these components have been used in the past to reproduce the large width of the IC SED peak (see e.g. Khatoon et al., 2024). Following Ghisellini & Tavecchio (2009); MAGIC Collaboration et al. (2019), we adopt erg s-1 and cm for the BLR, which leads to erg s-1 and cm using the typical relations between BLR, disc and torus (see Kaspi et al., 2007; Cleary et al., 2007), assuming in this case a covering factor for torus of , and a BLR reprocessing factor (based on the Hα line) . For the dusty torus, we assumed a typical temperature of 1000 K. We adopted a black hole mass of (Ghisellini et al., 2010). Finally, we considered the core located at a distance cm from the central engine and the blob at cm. We also take into account the possible absorption at GeV-TeV energies and the EBL absorption at VHE. We also evaluated the absorption due to the external radiation fields, which was found to be negligible.
Multi-zone models of blazar SEDs are characterised by a high degree of degeneracy, owing to the relatively large number of parameters. This degeneracy generally prevents the use of standard fit routines based on the minimisation of a cost function to effectively constrain the parameter space. In such cases, it is necessary to adopt a physically motivated scenario as guiding framework. In this work, we have investigated a scenario in which the evolution of the flare is driven by magnetic reconnection, which is strongly supported by the minute-scale variability detected at VHE with LST-1. In order to fit the SED of each BB, for each block we have used the parameters of the previous one as initial parameters, iterating until reaching a good representation of the SED. To reduce the number of free parameters, we have kept the magnetic fields , the emitting region sizes , the Lorentz factor and the Doppler factor fixed (with values in agreement with the limit values set by the variability analysis). Additionally, we minimise the changes of physical parameters in the core region between adjacent BBs to reflect the slow evolving nature of the core, unless the observed time interval between the centre of two adjacent BBs is longer than two days, the light crossing timescale of the core region in our reference frame. The blob, in contrast, is considered as the newly formed plasmoid from an unresolved magnetic reconnection layer in the core, whose parameters can vary considerably depending on the reconnection layer. Therefore, we have based our SED evolution in changes of the electron distribution of each region, that is, , , , , and . The parameter change was done in a way that is consistent with the flare’s evolution timescales. The resulting models for each BB are shown in Figs. 8 and 9, and the parameters describing each of them are summarised in Table 3. The consistency of these models with a magnetic reconnection scenario is discussed in Sect. 6.
The leptonic origin of the emission considered here is supported by recent results obtained by the IXPE (Imaging X-ray Polarimetry Explorer) satellite, based on the analysis of the broadband polarised emission of BL Lac (see Middei et al., 2023; Peirson et al., 2023). Under a leptonic framework where the X-ray spectrum is dominated by the high-energy IC scattering component, the X-ray polarisation degree is expected to be substantially lower than that of the optical band, which originates from electron synchrotron radiation (Liodakis et al., 2019; Peirson & Romani, 2019; Zhang et al., 2024). Middei et al. (2023) measured an upper limit of % for two IXPE observations of BL Lac during May and July 2022, consistent with this scenario given the optical polarisation degree.
Moreover, under strong flares where the soft X-ray emission has significant contribution from synchrotron radiation as observed from the much flatter X-ray spectrum — which is also the present case, see for instance BB3 with respect to e.g. BB8 when the source was at its brightest — one expects a significantly higher polarisation degree of the soft X-ray photons than that at hard X-ray energies, and even higher than the optical polarisation degree (Liodakis et al., 2019; Peirson & Romani, 2019; Zhang et al., 2024). A third IXPE observation of BL Lac, performed two days after the last LST-1 observation (November 27 to 29) resulted in a measured X-ray polarisation degree of over the entire X-ray band. After splitting the analysis in two bins accounting for the soft and hard X-ray regimes, almost all the polarised emission was found to be in the former, with a polarisation degree of . The synchrotron emission was estimated to comprise 44% of the total flux at the soft X-ray band (Peirson et al., 2023). After accounting only for the synchrotron emission and assuming that the 56% of the flux remaining comes from unpolarised SSC, the X-ray polarisation fraction was calculated to be 49%, much higher than that in optical frequencies (13%), and well within the expectations for a SSC scenario (Peirson & Romani, 2019). Therefore, the leptonic SED modelling performed here is consistent with the expected nature of the broadband emission based on both the spectral and the multi-wavelength polarisation analyses.
A last IXPE campaign on BL Lac during November 2023 resulted in an extremely high optical polarisation of %, not accompanied however with a high X-ray polarisation, leading to a 3 upper limit of 7.4%. This was observed during a period in which BL Lac was behaving as an LSP. Therefore, the X-ray emission was located in the high-energy SED component. These results strongly favour SSC multi-zone or energy-stratified models (see Zhang et al., 2024; Agudo et al., 2025), leaving little room for hadronic contribution, only in the framework of a hybrid model as shown by a detailed analysis of the broadband emission by Liodakis et al. (2025). However, as presented by these authors, hybrid models expect a hardening of the VHE spectrum that is not observed here in order to reproduce the multi-wavelength polarisation signatures.
6 Discussion
6.1 -ray and X-ray spectral behaviour
X-ray spectral trends were also observed in BL Lac prior to the 2022 high state (e.g. Weaver et al., 2020). Through an extensive SED modelling, Sahakyan & Giommi (2022) attributed the synchrotron component emerging in X-rays during periods of high state to a second emission zone located beyond the BLR, different from the one dominating the emission in the optical-UV band. The emission from this second region can produce a shift of the synchrotron peak towards higher frequencies, reaching values of even 10 Hz, well within the high-synchrotron-peaked (HSP) range. Indeed, BL Lac is well known for transitioning between the LSP regime — mainly during low flux states — to an IBL/HBL behaviour during active phases (see e.g. Ackermann et al., 2011; Nilsson et al., 2018). Starting on approximately 2019, the source has been going through several years of very high activity, culminating with its three brightest flares ever recorded in 2021, 2022 and 2024. Therefore, the X-ray spectral variability trends observed in the dataset analysed in this work can be related to the emergence of synchrotron emission from freshly accelerated particles in a secondary emission region along the jet. The shift of the X-ray emission from the high- to the low-energy peak of the SED can also be the reason behind the lower correlation of the hard X-ray and optical flux compared to the soft X-ray one, as reported in Appendix D.2. Furthermore, in our model the emission from the secondary region (blob) dominates also in the HE and VHE -ray bands. Therefore, the same freshly accelerated particles are also responsible for the intense flaring activity detected with Fermi-LAT and LST-1 via SSC, as well as for the observed brightening and shifting of the -ray spectrum towards higher energies (see Sect. 4).
6.2 Week-scale evolution of the broadband emission
We successfully model the broadband emission and evolution of the flare as a SSC two-zone model considering as well the EC contribution. The flare begins with emission dominated by the slowly evolving core, extending up to X-ray energies. Plasmoids (blobs) can develop within some considerably magnetised current sheets within the core through magnetic reconnection. These blobs become primarily responsible for the gamma-ray emission. At this early stage of the flare, the EC scattering also contributes to the gamma-ray emission. As magnetic reconnection sets in, particles in the blob are accelerated to higher energies, increasing its contribution also in the X-ray band. Our spectral fitting reveals several major ejections of plasmoids/blobs in the emission region, namely, BB4, BB6, BB8, and BB10. During all these episodes, a freshly produced blob dominates the gamma-ray emission, and can also have strong contributions in other wavelengths depending on the blob parameters. All those episodes are followed by a relatively low flux episode, where the newly formed blob cools and loses its energy. Since each blob is produced independently, their parameters are expected to be considerably different. On the other hand, the core evolves slowly, thus its physical parameters do not change much between different episodes.
This layout is in agreement with the correlation analysis performed in Appendix D.2, where the high degree of correlation between different bands can be attributed to the co-spatial nature of the broadband emission, with the exception of the lower correlation found with the radio and VHE bands. For the former, it is known that the inner regions of the jet are highly opaque to radio emission, that is typically assumed to come from an outer part (Max-Moerbeck et al., 2014). Several studies have in fact found long delays of hundreds of days between the radio and multi-wavelength emission of blazars (see e.g. Liodakis et al., 2018). We cannot test this hypothesis here as we focus on a shorter period. Nevertheless, such behaviour could explain the lower correlation of this band. We note, however, that the radio band as a lower observing cadence than that in optical, that can also be biasing the measured correlation. Regarding the latter band, the lower correlation of the VHE emission can be explained by the shorter cooling timescales of the highest-energy electrons responsible for the TeV emission. However, as for the radio band correlation, the lower observing cadence of the VHE data can introduce systematic effects in the correlation analysis, diluting the correlation of the TeV emission.
With the configuration assumed in Sect. 5, the low-energy photons from dusty torus are the ones dominating the EC contribution. This EC emission has been used in the past to model the broadband emission of BL Lac (Khatoon et al., 2024). In the present flare, owing to the very bright state in which the source was observed during most of the period, we find that the EC emission becomes relevant only at the early stages of the outburst, when the source is faintest (see the model for BB1 in Fig. 8). We have also considered the possible emission arising from the interplay between the core and blob regions, this is, low-energy photons from one region that can be IC-scattered by electrons of the other region. However, this was found to be negligible due to the compactness of the blob as a result of the size constraints derived from the VHE minute-scale variability. The absorption was also evaluated and taken into account, but was found to be negligible up to energies TeV, where the sources is not detected significantly and only upper limits are estimated.
Overall, the parameters are consistent with the typical values derived from a large sample of blazars (Ghisellini et al., 2010). The core region is mainly responsible for the emission from radio to UV wavelengths, and contributes significantly to the hard X-ray emission during those periods in which the X-ray emission is found in the transition from the low- to the high-energy peak of the SED. During the early moments of the flare, when the soft X-rays are also found in the IC peak, it dominates both the soft and hard X-ray emission. Changes in the parameters between adjacent BBs are rather small considering the time interval between consecutive blocks is generally longer than the light crossing time of the core, reflecting its slow evolution. On the other hand, during those periods when BL Lac fully transitions from LSP to ISP and the X-ray emission is entirely produced by synchrotron radiation, the low energy emission of the blob is the one becoming dominant in the X-ray band. We note that the values of the blob are on the high end of normal values. Zech & Lemoine (2021) interpreted these high values in other BL Lac objects as a thermal coupling of protons and electrons in a weakly magnetised shock, where protons transfer energy to electrons in the shock layer.
The main driver of the variability on day-to-week timescales is the evolution of the particle population of each region during the flare. Especially noticeable is the electron distribution behaviour of the blob during the brightest nights (BB4 and BB8), when it clearly reaches higher energies, in consistency with the IC peak shift towards higher energies reported in Sect. 4. Moreover, for these two nights, an injection of freshly accelerated particles was also needed to explain the remarkably high -ray emission. The blob is also characterised by a higher magnetisation and higher Doppler boosting than the core, with G, and G, , respectively. This, combined with the equipartition values of the core, close to equipartition but still 1, and its sometimes very hard spectral index, are consistent with a magnetic reconnection scenario (see e.g. Jormanainen et al., 2023), discussed in detail in Sect. 6.3. We note that the equipartition values of the blob are far from 1, with the electron energy density dominating on the order of magnitude of . This is a limitation that is regularly found in these models (see Tavecchio & Ghisellini, 2016). This might arise from assuming a still over-simplified picture of the relativistic jets while trying to capture all the observational features, including, week, day and minute scale variability. Nevertheless, the models presented here are able to accurately describe the evolution of the flare while still providing a reasonable solution to observational features such as multi-wavelength correlations, week-, day- and minute-scale VHE variability.
6.3 Origin of the minute-scale VHE variability
The very fast minute-scale variability demands a compact region with extreme particle acceleration (the blob in our fitting model) on top of the blazar zone (the core in our model) that governs the day-to-week-scale variability. Such a physical setup is often attributed to the mini-jet scenario under magnetic reconnection (Giannios et al., 2009; Petropoulou et al., 2016). A particularly high value of the blob Lorentz factor is also necessary to avoid the VHE -ray attenuation in the compact blob. The optical depth for VHE -rays is given by (Ackermann et al., 2016), where is the observed VHE photon energy, is the optical luminosity, which is the target photons for VHE -rays, and is the shortest observed variability timescale. We find that the Doppler factor can ensure the compact blob is optically thin to the observed VHE photons. This range of Doppler factor values is typically high with respect to those often observed in blazars. Hence, the mini-jet scenario is particularly relevant here to solve this problem and avoid the VHE -ray attenuation in the compact blob. Under the mini-jet scenario, the terminal bulk Lorentz factor of the plasmoid can reach approximately the magnetisation factor of the reconnection site in the co-moving frame (Sironi et al., 2016). As shown in Table 3, the magnetisation factor in the core, estimated by , is generally . The bulk Lorentz factor of the blob relative to the core can be approximated by (if they are both close to the line of sight), which is comparable to the magnetisation factor. Although our fitting model and the above estimates are rather simplified, they are consistent with the mini-jet in magnetic reconnection scenario (Nalewajko et al., 2011).
A plasmoid merger in magnetic reconnection can be an alternative scenario (Zhang et al., 2022). Previous particle-in-cell (PIC) simulations have shown that the SSC can exhibit much higher and faster variability than the synchrotron and EC, due to the highly concentrated electrons freshly accelerated at the plasmoid merging site. The magnetisation factor therein can drop below 1 even though the reconnection site has overall 1 magnetisation, consistent with the fitting parameters derived for the blob (Zhang et al., 2022). In principle, such collisions and mergers of plasma structures can happen in magnetised turbulence as well (Zhang et al., 2023). However, the very high intrinsic X-ray polarisation (49%, see Peirson et al., 2023) measured just two days after the last flare disfavours a turbulent environment.
In addition to the fast variability, the very hard particle spectrum in the core () also indicates a magnetic reconnection origin. Although we fit the radio to optical spectrum by the synchrotron emission from the core, usually the large-scale jet can have a considerable contribution to the radio emission. Therefore, the radio to optical emission intrinsic to the core in our model may require even harder particle spectrum. It is generally believed that only magnetic reconnection in a highly magnetised environment can result in particle spectrum harder than 2 (Guo et al., 2016, 2021). Interestingly, since the correlated quasi-periodic flux and polarisation variations detected by Jorstad et al. (2022) suggested kink instabilities in BL Lac, and the very high optical polarisation degree found in Agudo et al. (2025) indicates the relaxation of a magnetic spring in the jet, with the fast variability reported in the current paper, it is possible that the strong activities of BL Lac since roughly 2020 are associated with giant eruption of magnetised plasma from the central engine, likely due to the magnetically-arrested-disc jet scenario.
Under this picture, our two-zone model essentially consists of a segment of the kinked jet (the core) and plasmoids (the blob) produced via magnetic reconnection that is triggered by the kinked jet from time to time. Since they are co-spatial and physically connected, all bands above the synchrotron self absorption limit should have some correlations. However, the production of plasmoids and their mergers in reconnection is rather random as shown in numerical simulations (Sironi et al., 2016; Zhang et al., 2022). Hence the VHE emission that is fully dominated by the blob can appear less correlated to the optical emission that is fully dominated by the core. Moreover, the models presented here propose a simplified picture of the magnetic reconnection scenario. As the flare develops, plasmoids grow and can be ejected from the reconnection layer with various physical conditions as suggested by recent PIC simulations (Sironi et al., 2016; Zhang et al., 2024). While a fully numerical treatment of plasma dynamics in reconnection is far beyond the scope of this paper, we model these plasmoids as one blob with highly variable physical parameters as shown in many previous studies (e.g. Christie et al., 2019). We note that the two-zone model has significant parameter degeneracy, thus our fitting parameters may not be unique.
Additionally, we have calculated the blob synchrotron and SSC cooling timescales in the observer frame for the electrons responsible for the GeV-TeV -ray emission. The synchrotron cooling timescales were estimated as (see e.g. Abe et al., 2025)
where is the electron mass, is the Thomson cross-section, is the magnetic field of the blob and is the Lorentz factor of the electrons. The blob SSC cooling timescales in the observer frame were computed as , with expressed in the blob reference frame as (Moderski et al., 2005)
where is the photon energy in the rest frame of the blob in units of the electron rest mass energy, , being the photon frequency in the blob frame, is defined by Eq. (C5) in Moderski et al. (2005), is the synchrotron photon number density in the blob rest frame, computed as in Abe et al. (2025). The blob SSC cooling timescales were found to be shorter than the synchrotron ones, consistent with the Compton dominance of – shown by the blob SED. In Table 3 we report the blob SSC cooling timescales at . For the two nights with intranight variability, our models predict SSC cooling timescales of the order of few minutes for electrons radiating at GeV-TeV, in agreement with the minute scale variability observed. Therefore, with all the considerations mentioned above, one possible scenario is that the observed flares at VHE rays are produced by freshly accelerated particles via magnetic reconnection with a cooling dominated by SSC cooling.
We note that the scenario proposed here is based on a magnetically-dominated region. However, an alternative picture based on a two-zone model with a weakly magnetised plasma can still offer a good fitting to the SED as long as the blob has a higher Lorentz factor than that of the core, and larger than the expected minimum Lorentz factor obtained from the TeV fast variability, estimated above. The most natural way to explain higher Lorentz factor in a small blob compared to the overall blazar zone is magnetic reconnection. Nevertheless, and owing to the high degeneracy of blazar SED models (as commented in Sect. 5), other combinations of parameters can lead to valid representations of the SED and its evolution, possibly relying on non-magnetically-dominated scenarios.
7 Conclusions
BL Lac has shown one of its three brightest -ray and multi-wavelength flares ever between September and November 2023. We have characterised its HE and VHE emission, extending up to a few TeV during the brightest periods of the flare. One of the key features of this event is the significant remarkably fast VHE -ray flux variability, with doubling flux timescales as short as 8.3 minutes, which sets tight constraints to the size of the VHE emitting region. This is also among the fastest flux variations ever detected at VHEs from blazars.
The HE-VHE spectrum shows a shift of the peak towards higher energies at the highest stages of the flare. Accompanying this behaviour we also report a softer-when-brighter evolution of the X-ray spectrum due to a clear transition of the X-ray emission from IC-dominated to synchrotron-dominated, meaning that the source also transitions from LSP to ISP, as seen in other periods for BL Lac in the past (Nilsson et al., 2018).
The broadband emission was modelled within a leptonic scenario, motivated by recent results reported by IXPE (Middei et al., 2023; Peirson et al., 2023; Agudo et al., 2025; Liodakis et al., 2025). One-zone models were not successful when trying to explain the broadband emission. Therefore, a co-spatial two-zone SSC model incorporating also the dusty torus EC contribution was used. Within this scenario, the evolution of the jet is driven by changes on the electron population of both regions, with injection of freshly accelerated electrons being the cause of the brightest outbursts. The minute-scale variability observed constrains the size of the compact blob region to be cm for the adopted Doppler factor , consistent with the value used here. This blob region is therefore the responsible of the intranight -ray variability, that can be associated with a plasmoid under a magnetic reconnection event.
Appendix A Night-wise LST-1 data analysis results
In Table 4 we compile and summarise the night-wise results of the LST-1 data analysis of BL Lac, that is, the effective observing time after data quality selection, the significance, and the flux of the source above 200 GeV.
| Day | Observation | Significance | Fluxa | Day | Observation | Significance | Fluxa | |
| Time [h] | [] | [ cm-2 s-1] | Time [h] | [] | [ cm-2 s-1] | |||
| 21/09 | 1.71 | 0.32 | 27/10 | 1.36 | 6.48 | |||
| 01/10 | 1.65 | 1.22 | 29/10 | 0.72 | 6.19 | |||
| 03/10 | 2.86 | 2.28 | 31/10 | 1.73 | 5.08 | |||
| 15/10 | 1.31 | 7.55 | 01/11 | 1.35 | 5.22 | |||
| 17/10 | 2.14 | 10.77 | 13/11 | 1.51 | 30.47 | |||
| 18/10 | 2.48 | 1.14 | 15/11 | 0.76 | 2.20 | |||
| 19/10 | 2.67 | 3.26 | 16/11 | 0.88 | 0.54 | |||
| 20/10 | 2.50 | 48.73 | 17/11 | 0.87 | 3.01 | |||
| 21/10 | 1.87 | 0.70 | 18/11 | 1.01 | 0.68 | |||
| 23/10 | 1.41 | 1.70 | 19/11 | 0.57 | 0.66 | |||
| 25/10 | 2.25 | 0.13 | 25/11 | 0.76 | 10.53 | |||
| 26/10 | 1.23 | 4.67 |
Appendix B X-ray data results
In Table 5 we report the summary of the results of the Swift-XRT analysis, presented in Sect. 2.3. For the observations in which the log-parabola model was preferred over the power-law one with a 3 significance, Table 6 reports the best-fit parameters of the log-parabola model. The logpar implementation in XSPEC was employed in the fit procedures. The derived SEDs are represented in Fig. 5, along with those derived from the UVOT analysis in the optical-UV range.
| Day | MJD | ObsID | Exposure | Significance [] | Flux | Flux | Spectral index |
| [s] | Log-parabola model | [10] | [10] | ||||
| 09/16 | 59838.6 | 00014925014 | 1981 | 1.7 | |||
| 09/18 | 59840.5 | 00014925015 | 1994 | 1.9 | |||
| 09/20 | 59842.3 | 00014925016 | 572 | 0.9 | |||
| 09/23 | 59845.5 | 00014925017 | 705 | 0.3 | |||
| 10/17 | 59869.8 | 00014925018 | 1102 | 1.5 | |||
| 10/17 | 59869.9 | 00014925019 | 1412 | 0.4 | |||
| 10/22 | 59874.4 | 00014925021 | 1978 | 3.7 | |||
| 10/23 | 59875.8 | 00014925020 | 1832 | 4.3 | |||
| 10/25 | 59877.0 | 00014925023 | 1008 | 3.8 | |||
| 10/26 | 59878.3 | 00096990012 | 1034 | 1.9 | |||
| 10/29 | 59881.1 | 00014925024 | 1661 | 3.4 | |||
| 11/01 | 59884.1 | 00014925026 | 951 | 1.1 | |||
| 11/04 | 59887.3 | 00014925027 | 1240 | 2.3 | |||
| 11/07 | 59890.1 | 00014925028 | 1035 | 4.2 | |||
| 11/12 | 59895.4 | 00096990013 | 849 | 1.9 | |||
| 11/13 | 59896.3 | 00014925030 | 1278 | 2.2 | |||
| 11/21 | 59904.8 | 00014925032 | 1681 | 2.4 | |||
| 11/23 | 59906.0 | 00014925033 | 1447 | 2.7 | |||
| 11/25 | 59908.4 | 00096990014 | 865 | 2.0 | |||
| 11/25 | 59908.7 | 00014925034 | 2021 | 2.5 | |||
| 11/26 | 59909.5 | 00096990015 | 905 | 2.6 | |||
| 11/27 | 59910.2 | 00096990016 | 834 | 2.8 | |||
| 11/28 | 59911.9 | 00089562001 | 1473 | 3.7 | |||
| 11/29 | 59912.2 | 00096990017 | 826 | 2.7 | |||
| 11/30 | 59913.4 | 00096990018 | 895 | 3.7 | |||
| 12/01 | 59914.1 | 00096990019 | 960 | 3.2 |
| Day | MJD | ObsID | Significance [] | ||||
| Log-parabola model | [keV] | [10] | |||||
| 10/22 | 59874.4 | 00014925021 | 3.7 | 1 | |||
| 10/23 | 59875.8 | 00014925020 | 4.3 | 1 | |||
| 10/25 | 59877.0 | 00014925023 | 3.8 | 1 | |||
| 10/29 | 59881.1 | 00014925024 | 3.4 | 1 | |||
| 11/07 | 59890.1 | 00014925028 | 4.2 | 1 | |||
| 11/28 | 59911.9 | 00089562001 | 3.7 | 1 | |||
| 11/30 | 59913.4 | 00096990018 | 3.7 | 1 | |||
| 12/01 | 59914.1 | 00096990019 | 3.2 | 1 |
The right panel of Fig. 5 shows the scatter plot of the X-ray intrinsic power-law index and flux reconstructed from the 50 Swift-XRT observations of BL Lac from March 2022 to January 2023 (Sect. 2.3). A softer-when-brighter trend can be highlighted, that is, higher fluxes correspond to softer spectra. The correlation between and yields a Pearson coefficient of , with a -value under the null hypothesis of non-correlation of . Hence, a positive correlation is statistically significant. In addition, we observe that the X-ray photon index transitions from values 2.5, up to 1.4, indicating a spectrum dominated by the high-energy emission when , and indicating the emergence of an X-ray synchrotron component for . The transition is indeed visible in the SED models presented in Sect. 5 and Appendix F, as well as in the left panel of Fig. 5, where the X-ray spectrum is found in the increasing part of the high-energy SED peak during low emission states, whereas it corresponds to the decreasing part of the synchrotron peak for high-flux periods. For some of the observations, concave log-parabola models were preferred over power-law ones with a significance 3 to describe the intrinsic X-ray spectrum of BL Lac (Table 6). This indicates that the keV band hosts the valley resulting from the transition between the low-energy and high-energy radiative components (see left panel of Fig. 5).
Appendix C Swift-UVOT data results
In Table 7, we report the results of the analyses of Swift-UVOT data from the observations carried out from September to November 2022. For each observation, the flux densities in the Swift-UVOT photometric bands are reported, corrected for both Galactic extinction and host galaxy contribution (see Sect. 2.4 for more details). The optical-UV SEDs over this period are shown in the left panel of Fig. 5.
| Day | MJD | ObsID | Exposure | V | B | U | W1 | M2 | W2 |
| [s] | [mJy] | [mJy] | [mJy] | [mJy] | [mJy] | [mJy] | |||
| 09/16 | 59838.6 | 00014925014 | 1948 | ||||||
| 09/18 | 59840.5 | 00014925015 | 1898 | ||||||
| 09/20 | 59842.3 | 00014925016 | 547 | ||||||
| 09/23 | 59845.5 | 00014925017 | 680 | ||||||
| 10/17 | 59869.8 | 00014925018 | 1450 | - | - | - | - | - | |
| 10/17 | 59869.9 | 00014925019 | 1453 | - | - | - | - | - | |
| 10/22 | 59874.4 | 00014925021 | 1925 | ||||||
| 10/23 | 59875.8 | 00014925020 | 1758 | ||||||
| 10/25 | 59877.0 | 00014925023 | 961 | ||||||
| 10/26 | 59878.3 | 00096990012 | 1009 | ||||||
| 10/29 | 59881.1 | 00014925024 | 1610 | ||||||
| 11/01 | 59884.1 | 00014925026 | 928 | ||||||
| 11/04 | 59887.2 | 00014925027 | 1188 | ||||||
| 11/07 | 59890.1 | 00014925028 | 990 | ||||||
| 11/12 | 59895.4 | 00096990013 | 826 | ||||||
| 11/13 | 59896.3 | 00014925030 | 1254 | ||||||
| 11/21 | 59904.8 | 00014925032 | 1655 | ||||||
| 11/23 | 59906.0 | 00014925033 | 1397 | ||||||
| 11/25 | 59908.4 | 00096990014 | 839 | ||||||
| 11/25 | 59908.6 | 00014925034 | 1985 | ||||||
| 11/26 | 59909.5 | 00096990015 | 879 | ||||||
| 11/27 | 59910.2 | 00096990016 | 809 | ||||||
| 11/28 | 59911.9 | 00089562001 | 1446 | ||||||
| 11/29 | 59912.2 | 00096990017 | 801 | ||||||
| 11/30 | 59913.4 | 00096990018 | 870 | ||||||
| 12/01 | 59914.1 | 00096990019 | 914 |
Appendix D Multi-wavelength light curve analysis
D.1 Variability
The HE and VHE -ray flare studied here and shown in Fig. 1 corresponds to one of the three highest -ray activities from this source, together with the roughly equally bright outbursts that occurred on August 2021 and October 2024. The night-wise VHE variability ranges from non-detections during which the emission of the source was low, up to a flux of 1.9 Crab Units (C.U.), a factor 10 with respect to the average emission of 0.23 C.U. over the whole period (see Table 4 for the night-wise results). This variability becomes even more evident when evaluating the flux on a shorter timescale (see Sect. 3 for more details), reaching a maximum emission of 2.8 C.U. above 200 GeV for the night of November 13, and 4.4 C.U. above 100 GeV for October 20.
Extreme variability is also present in the HE emission of BL Lac, with several rises and decays of the flux over time. The LAT data show an extraordinarily bright period, with the flux changing almost a factor 70 from the minimum to the maximum emission detected, and reaching a flux of cm-2 s-1 on October 16, being amongst the brightest blazar flares ever detected by this instrument. The LST-1 light curve does not reveal a similar VHE flux increase around October 16, however we note that the LAT light curve is based on a 12-hour binning and the highest state occurs during day time for LST-1, therefore after the VHE -ray observations. Hence, we can not exclude the presence of a fast VHE flare such as those observed during October 20 and November 13, with a flux above 200 GeV that significantly increased after the LST-1’s observations.
As for the X-ray flux between 0.3 and 10 keV, we observe an energy flux variation of a factor 11, similar to that in the VHE band, ranging from 22 to 2, in units of erg cm-2 s-1. However, we stress that the observing coverage is relatively low especially for the first half of the campaign. Nevertheless, during the second half, the X-ray emission seems to follow the same variability pattern as that seen in HE rays. Remarkably, the highest X-ray emission is detected during a period in which the HE -ray emission is moderate. Unfortunately, no simultaneous LST-1 observation is available to compare the most energetic -ray emission with the high-energy end of the synchrotron component, located in the X-ray band for this observation, see Sect. 5 and Appendix C.
Large variability is also observed in the optical-UV bands. Taking the UVOT UV band as reference, we measure an average emission of 8.5 mJy, varying a factor 20 between 0.9 and 18.8 mJy. The same increase is measured in the optical band, where the flux density goes from 8.6 mJy at the lowest state, up to 192.9 mJy at the peak of the flare. Finally, as commonly observed for -ray blazars, the radio bands show the lowest variability across all bands. At a frequency of 37 GHz, the emission changes from 5.47 and 7.83 Jy, only a factor 1.4, contrasting with the large variability observed at other wavelengths.
D.2 Correlations
As shown by Fig. 1, the multi-wavelength emission seems to follow similar variability patterns in several bands. In order to quantify the possible correlations between the different bands, we have computed the Spearman’s linear correlation coefficient for each wavelength with respect to the band. We use strictly simultaneous data within 0.1 days to avoid biasing the results as much as possible, with the exception of LAT data, for which we consider the -band observations within each 12-hour bin.
No significant correlation is observed between the VHE and optical light curves, for which we measure a coefficient (-value ). This contrasts with a mild correlation between the HE and VHE -ray bands, that show (-value ). A very high correlation degree is observed between the optical and UV bands, for which we obtain coefficients and (-values and ), respectively (using the and bands in each case). A high correlation is also observed between the optical and HE -ray emission, with (-value ). The X-ray emission on the other hand shows a slightly smaller value, indicating a hint of correlation with the optical emission. The estimated coefficient for this pair of light curves is (). We have evaluated the cause of this somewhat lower correlation with respect to the optical-UV bands by calculating the correlation coefficient for the soft (0.3-2 keV) and hard (2-10 keV) X-ray bands. The former shows again a very high degree of correlation with the band (, ), whereas the latter shows a lower correlation degree (, ). This can be attributed to the transition of the X-ray emission from the high-energy to the synchrotron SED bump. The hard X-ray band hosts more frequently the valley between both peaks compared to the soft X-ray band (see Appendix C), hence lowering the level of correlation with the optical emission. Finally, the radio emission shows the lowest correlation degree, with (). This is not surprising since, as reported in Appendix D.1, the radio emission is the least variable, therefore not following exactly the same variability features as the optical band.
Appendix E Joint HE-VHE SED fits
In Figs. 6 and 7 we show the results of the joint LAT-LST-1 spectral characterisation described in Sect. 4 for all the BBs defined in our analysis. These SEDs correspond to the spectra whose parameters are summarised in Table 2. The spectral shape of the joint analysis at VHE -ray energies is also compared with that obtained from an LST-only analysis.
Appendix F SED models
This appendix gathers the SED models for all 10 BBs defined in Table 2. The fits are represented in Figs. 8 and 9, and correspond to the models defined by the parameters reported in Table 3.
Acknowledgements
Acknowledgements: We sincerely thank the anonymous referee for their thorough review of the manuscript and helpful comments.
We gratefully acknowledge financial support from the following agencies and organisations:
Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) Grant 309053/2022-6 and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) Grants E-26/200.532/2023 and E-26/211.342/2021, Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), Fundação de Apoio à Ciência, Tecnologia e Inovação do Paraná - Fundação Araucária, Ministry of Science, Technology, Innovations and Communications (MCTIC), Brasil;
Ministry of Education and Science, National RI Roadmap Project DO1-153/28.08.2018, Bulgaria;
Croatian Science Foundation (HrZZ) Project IP-2022-10-4595, Rudjer Boskovic Institute, University of Osijek, University of Rijeka, University of Split, Faculty of Electrical Engineering, Mechanical Engineering and Naval Architecture, University of Zagreb, Faculty of Electrical Engineering and Computing, Croatia;
Ministry of Education, Youth and Sports, MEYS LM2023047, EU/MEYS CZ.02.1.01/0.0/0.0/16_013/0001403, CZ.02.1.01/0.0/0.0/18_046/0016007, CZ.02.1.01/0.0/0.0/16_019/0000754, CZ.02.01.01/00/22_008/0004632 and CZ.02.01.01/00/23_015/0008197 Czech Republic;
CNRS-IN2P3, the French Programme d’investissements d’avenir and the Enigmass Labex,
This work has been done thanks to the facilities offered by the Univ. Savoie Mont Blanc - CNRS/IN2P3 MUST computing center, France;
Max Planck Society, German Bundesministerium für Forschung, Technologie und Raumfahrt (Verbundforschung / ErUM), the Deutsche Forschungsgemeinschaft (SFB 1491) and the Lamarr-Institute for Machine Learning and Artificial Intelligence, Germany;
Istituto Nazionale di Astrofisica (INAF), Istituto Nazionale di Fisica Nucleare (INFN), Italian Ministry for University and Research (MUR), and the financial support from the European Union – Next Generation EU under the project IR0000012 - CTA+ (CUP C53C22000430006), announcement N.3264 on 28/12/2021: “Rafforzamento e creazione di IR nell’ambito del Piano Nazionale di Ripresa e Resilienza (PNRR)”;
ICRR, University of Tokyo, JSPS, MEXT, Japan;
JST SPRING - JPMJSP2108;
Narodowe Centrum Nauki, grant number 2023/50/A/ST9/00254, Poland;
The Spanish groups acknowledge the Spanish Ministry of Science and Innovation and the Spanish Research State Agency (AEI) through the government budget lines
PGE2022/28.06.000X.711.04,
28.06.000X.411.01 and 28.06.000X.711.04 of PGE 2023, 2024 and 2025,
and grants PID2019-104114RB-C31, PID2019-107847RB-C44, PID2019-105510GB-C31, PID2019-104114RB-C33, PID2019-107847RB-C43, PID2019-107847RB-C42, PID2019-107988GB-C22, PID2021-124581OB-I00, PID2021-125331NB-I00, PID2022-136828NB-C41, PID2022-137810NB-C22, PID2022-138172NB-C41, PID2022-138172NB-C42, PID2022-138172NB-C43, PID2022-139117NB-C41, PID2022-139117NB-C42, PID2022-139117NB-C43, PID2022-139117NB-C44, PID2022-136828NB-C42, PID2024-155316NB-I00, PDC2023-145839-I00 funded by the Spanish MCIN/AEI/10.13039/501100011033 and by ERDF/EU and NextGenerationEU PRTR; CSIC PIE 202350E189; the ”Centro de Excelencia Severo Ochoa” program through grants no. CEX2020-001007-S, CEX2021-001131-S, CEX2024-001442-S; the ”Unidad de Excelencia María de Maeztu” program through grants no. CEX2019-000918-M, CEX2020-001058-M; the ”Ramón y Cajal” program through grants RYC2021-032991-I funded by MICIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR and RYC2020-028639-I; the ”Juan de la Cierva-Incorporación” program through grant no. IJC2019-040315-I and ”Juan de la Cierva-formación”’ through grant JDC2022-049705-I; the “Viera y Clavijo” postdoctoral program of Universidad de La Laguna, funded by the Agencia Canaria de Investigación, Innovación y Sociedad de la Información. They also acknowledge the ”Atracción de Talento” program of Comunidad de Madrid through grant no. 2019-T2/TIC-12900; “MAD4SPACE: Desarrollo de tecnologías habilitadoras para estudios del espacio en la Comunidad de Madrid” (TEC-2024/TEC-182) project, Doctorado Industrial (IND2024/TIC34250) and Ayudas para la contratación de personal investigador predoctoral en formación (PIPF-2023/TEC-29694) funded by Comunidad de Madrid; the La Caixa Banking Foundation, grant no. LCF/BQ/PI21/11830030; Junta de Andalucía under Plan Complementario de I+D+I (Ref. AST22_0001) and Plan Andaluz de Investigación, Desarrollo e Innovación as research group FQM-322; Project ref. AST22_00001_9 with funding from NextGenerationEU funds; the “Ministerio de Ciencia, Innovación y Universidades” and its “Plan de Recuperación, Transformación y Resiliencia”; “Consejería de Universidad, Investigación e Innovación” of the regional government of Andalucía and “Consejo Superior de Investigaciones Científicas”, Grant CNS2023-144504 funded by MICIU/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR, the European Union’s Recovery and Resilience Facility-Next Generation, in the framework of the General Invitation of the Spanish Government’s public business entity Red.es to participate in talent attraction and retention programmes within Investment 4 of Component 19 of the Recovery, Transformation and Resilience Plan; Junta de Andalucía under Plan Complementario de I+D+I (Ref. AST22_00001), Plan Andaluz de Investigación, Desarrollo e Innovación (Ref. FQM-322). “Programa Operativo de Crecimiento Inteligente” FEDER 2014-2020 (Ref. ESFRI-2017-IAC-12), Ministerio de Ciencia e Innovación, 15% co-financed by Consejería de Economía, Industria, Comercio y Conocimiento del Gobierno de Canarias; the ”CERCA” program and the grants 2021SGR00426 and 2021SGR00679, all funded by the Generalitat de Catalunya; and the European Union’s NextGenerationEU (PRTR-C17.I1). This work is funded/Co-funded by the European Union (ERC, MicroStars, 101076533). This research used the computing and storage resources provided by the Port d’Informació Científica (PIC) data center.
State Secretariat for Education, Research and Innovation (SERI) and Swiss National Science Foundation (SNSF), Switzerland;
The research leading to these results has received funding from the European Union’s Seventh Framework Programme (FP7/2007-2013) under grant agreements No 262053 and No 317446;
This project is receiving funding from the European Union’s Horizon 2020 research and innovation programs under agreement No 676134;
ESCAPE - The European Science Cluster of Astronomy & Particle Physics ESFRI Research Infrastructures has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement no. 824064.
H.Z. is supported by NASA under award number 80GSFC24M0006.
Some of the data are based on observations collected at the Observatorio de Sierra Nevada; which is owned and operated by the Instituto de Astrofísica de Andalucía (IAA-CSIC).
This article is based on observations made with the IAC80 operated on the island of Tenerife by the Instituto de Astrofísica de Canarias in the Spanish Observatorio del Teide.
This work makes use of observations from the Las Cumbres Observatory global telescope network.
This work makes use of data from the All-Sky Automated Survey for Supernovae (ASAS-SN).
This work has made use of data from the Joan Oró Telescope (TJO) of the Montsec Observatory (OdM), which is owned by the Catalan Government and operated by the Institute for Space Studies of Catalonia (IEEC).
Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Oskar Klein Center at Stockholm University, the University of Maryland, University of California, Berkeley , the University of Wisconsin at Milwaukee, University of Warwick, Ruhr University, Cornell University, Northwestern University and Drexel University. Operations are conducted by COO, IPAC, and UW.
This publication makes use of data obtained at the Metsähovi Radio Observatory, operated by the Aalto University.
The Submillimeter Array is a joint project between the Smithsonian Astrophysical Observatory and the Academia Sinica Institute of Astronomy and Astrophysics and is funded by the Smithsonian Institution and the Academia Sinica. Maunakea, the location of the SMA, is a culturally important site for the indigenous Hawaiian people; we are privileged to study the cosmos from its summit.
Author contributions: D. Cerasole: Swift-XRT/UVOT data analysis, SED modelling and discussion, paper edition. G. Emery: LST-1 data analysis, systematic evaluation, paper edition. D. Morcuende: LST-1 data analysis, paper edition. J. Otero-Santos: project coordination, LST-1, Fermi-LAT and multi-wavelength data analysis, SED modelling and interpretation, paper edition. H. Zhang: theoretical interpretation, paper edition. The rest of the authors have contributed in one or several of the following ways: design, construction, maintenance and operation of the instrument(s) used to acquire the data; preparation and/or evaluation of the observation proposals; data acquisition, processing, calibration and/or reduction; production of analysis tools and/or related Monte Carlo simulations; discussion and approval of the contents of the draft.

### 07 Apr 2026 v3

@@
-VHE gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022
+Very-high-energy gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022
 Abstract
 Context.
-BL Lacertae (BL Lac), the archetype blazar of its subclass and one of the most studied blazars in the last decades, has gone through a series of major multi-wavelength outbursts since 2020, resulting in its highest recorded -ray flare up to date between September and November 2022 together with those from August 2021 and October 2024.
+BL Lacertae (BL Lac), the archetypal blazar of its subclass and one of the most studied blazars of recent decades, has undergone a series of major multi-wavelength outbursts since 2020, resulting in its highest recorded -ray flare to date between September and November 2022, together with those from August 2021 and October 2024.
 Aims.
-We characterise the -ray and multi-wavelength emission and spectral energy distribution (SED) of BL Lac, as well as their evolution during the major and extended -ray and multi-wavelength flare occurring between September and November 2022.
+We characterised the -ray and multi-wavelength emission and spectral energy distribution (SED) of BL Lac, as well as their evolution during the major and extended -ray and multi-wavelength flare that occurred between September and November 2022.
 Methods.
-We evaluate the variability of the flare, with focus on the nights of October 20 and November 13, when clear intranight very-high-energy (VHE, GeV) -ray variability is observed.
-We model the -ray and broadband SEDs during periods of stable emission identified with a Bayesian block analysis, interpreting their evolution of the flare from the variability of the relativistic particles and physical parameters of the jet.
+We evaluated the variability of the flare, focusing on the nights of October 20 and November 13, 2022 when clear intranight very-high-energy (VHE, GeV) -ray variability was observed.
+We modelled the -ray and broadband SEDs during periods of stable emission identified with a Bayesian block analysis and interpreted the flare’s evolution in terms of the variability in the relativistic particles and the jet’s physical parameters.
 Results.
-The VHE emission shows an average flux of 0.23 Crab Units (C.U.) above 200 GeV during this flare and a variability amplitude of more than a factor 10.
-Intranight doubling-flux variations as fast as 8 minutes are observed during the nights of October 20 and November 13, when maximum fluxes of 4.4 C.U.
+During this flare, the VHE emission shows an average flux of 0.23 Crab units (C.U.) above 200 GeV and a variability amplitude of more than a factor of ten.
+We observe intranight flux-doubling variations as fast as 8 minutes during the nights of October 20 and November 13, 2022 with maximum fluxes of 4.4 C.U.
 above 100 GeV and 2.8 C.U.
-above 200 GeV are reached.
-The spectral analysis reveals a transition of the X-ray emission from the high- to the low-energy SED peak, and a shift of the -ray peak towards higher energies.
-The broadband emission was interpreted within a leptonic two-zone model in which intranight variability is explained as magnetic reconnection in a compact region closely oriented with the line of sight, while weekly-scale variations can be explained as variations of the relativistic electron distributions and the injection of freshly accelerated particles.
+above 200 GeV.
+The spectral analysis reveals a transition of the X-ray emission from the high- to the low-energy SED peak and a shift in the -ray peak towards higher energies.
+We interpret the broadband emission within a leptonic two-zone model in which intranight variability is explained as magnetic reconnection in a compact region closely orientated with the line of sight, while variations in the relativistic electron distributions and the injection of freshly accelerated particles explain the weekly scale variations.
 Key Words.:
 Galaxies: active – BL Lacertae objects: general – BL Lacertae objects: individual: BL Lacertae – Galaxies: jets – Galaxies: nuclei – gamma rays: galaxies1 Introduction
-Relativistic jets launched from blazars — active galactic nuclei (AGN) with their jets oriented towards our line of sight — are known to be extremely efficient and natural particle accelerators, which are able to emit in all the electromagnetic spectrum.
-In fact, they are one of the few sources that can emit up to TeV energies.
+Relativistic jets launched from blazars, which are active galactic nuclei (AGNs) with their jets orientated towards our line of sight, are extremely efficient and natural particle accelerators that emit across the entire the electromagnetic spectrum.
+In fact, they are among the few sources that can emit up to TeV energies.
 Blazars are also highly variable on all timescales.
-Depending on the properties of their optical spectrum (absence/presence of broad emission lines, see Stickel et al., 1991), they can be classified into two main sub-classes: BL Lac type objects and flat spectrum radio quasars (FSRQs) (Urry & Padovani, 1995).
+Depending on the properties of their optical spectrum (absence or presence of broad emission lines; see Stickel et al., 1991), they can be classified into two main sub-classes: BL Lac type objects and flat spectrum radio quasars (FSRQs) (Urry and Padovani, 1995).
 The broadband spectral energy distribution (SED) of blazars, extending from radio to -ray energies, has a characteristic double-peak shape.
 The low-energy component is produced by synchrotron radiation (see Konigl, 1981).
-The invoked mechanisms behind the high-energy component fall in two groups.
-An interpretation based on leptonic processes explains this emission through Inverse Compton (IC) scattering with low-energy photons, either synchrotron radiation via synchrotron self Compton (SSC, see Konigl, 1981; Maraschi et al., 1992) scattering, or with radiation from outside the jet via External Compton (EC, see Dermer & Schlickeiser, 1994) scattering.
-However, interpretations based on a hadronic nature are also used in some cases, relying on processes like proton synchrotron radiation or photo-pion production to explain the high-energy emission of blazars (see e.g.
-Mannheim & Biermann, 1992; Aharonian, 2000).
-BL Lacertae (hereafter BL Lac), the archetype of the BL Lac blazar subclass located at redshift (Miller et al., 1978), has historically shown remarkable variability with several episodes of minute- to hour-scale variations (see for instance Abeysekara et al., 2018; MAGIC Collaboration et al., 2019; Jorstad et al., 2022), becoming one of the most monitored and studied blazars over the years.
+The mechanisms invoked to explain the high-energy component fall into two groups.
+An interpretation based on leptonic processes explains this emission through inverse Compton (IC) scattering with low-energy photons, either from synchrotron radiation via synchrotron self Compton (SSC; see Konigl, 1981; Maraschi et al., 1992) scattering or from radiation from outside the jet via external Compton (EC; see Dermer and Schlickeiser, 1994) scattering.
+However, hadronic-based interpretations are also used in some cases and rely on processes such as proton synchrotron radiation or photo-pion production to explain the high-energy emission of blazars (see, e.g.
+Mannheim and Biermann, 1992; Aharonian, 2000).
+BL Lacertae (hereafter BL Lac), the archetype of the BL Lac blazar subclass located at redshift (Miller et al., 1978), historically shows remarkable variability with several episodes of minute- to hour-scale variations (see, for instance Abeysekara et al., 2018; MAGIC Collaboration et al., 2019; Jorstad et al., 2022), making it one of the most monitored and studied blazars to date.
 It is classified as an intermediate-synchrotron-peaked (ISP) blazar according to the peak frequency of its synchrotron emission ().
-However, it is known for transitioning between the low-synchrotron-peaked (LSP, ) and ISP types, depending on its activity (see e.g.
+However, it is known to transition between low-synchrotron-peaked (LSP; ) and ISP states, depending on its activity (see, e.g.
 Ackermann et al., 2011; Nilsson et al., 2018).
-It has been detected several times in the very-high-energy (VHE, GeV) -ray band, always during multi-wavelength flaring emission states (Albert et al., 2007; Arlen et al., 2013).
-These periods of variability on very short timescales imply very compact emitting regions involving the most energetic particle acceleration processes, and challenging in some cases the theoretical models used for interpreting the broadband emission of blazars (Henri & Saugé, 2006).
-Here we report the results of an extreme flare observed from BL Lac by the Large-Sized Telescope prototype (LST-1), taking place between September and November 2022, and triggered after the detection of enhanced activity by the Fermi satellite (La Mura, 2022).
-This flare is one of the three brightest events ever detected in VHEs from BL Lac, together with similar events occurring in 2021 and 2024, and it is also the most long-lasting flaring period ever recorded for this source in the VHE domain, lasting for about 3 months, with significant variations in day, hour and minute timescales.
-This remarkable flare was also covered by several multi-wavelength facilities, providing a rich dataset for interpreting the multi-wavelength emission and variability of BL Lac in this period (see Sect.
+It has been detected several times in the very-high-energy (VHE; GeV) -ray band, always during multi-wavelength flaring emission states (Albert et al., 2007; Arlen et al., 2013).
+These periods of variability on very short timescales imply very compact emitting regions involving the most energetic particle acceleration processes, which in some cases challenge the theoretical models used to interpret the broadband emission of blazars (Henri and Saugé, 2006).
+Here we report the results of an extreme flare observed from BL Lac by the Large-Sized Telescope prototype (LST-1), which took place between September and November 2022 and was triggered after the detection of enhanced activity by the Fermi satellite (La Mura, 2022).
+This flare is one of the three brightest events ever detected in VHEs from BL Lac, together with similar events that occurred in 2021 and 2024.
+It is also the longest flaring period ever recorded for this source in the VHE domain, lasting about three months, with significant variations on day-, hour-, and minute-timescales.
+Several multi-wavelength facilities also covered this remarkable flare, providing a rich dataset for interpreting the multi-wavelength emission and variability of BL Lac during this period (see Sect.
 2).
@@
 We characterise the minute-scale -ray variability observed during this period (Sect.
-3), that sets important constrains to the nature of the emitting region.
-The spectral behaviour of the source is evaluated (Sect.
-4), and its broadband emission is modelled within AGN radiative models (Sect.
+3), which sets important constraints on the nature of the emitting region.
+We evaluate the spectral behaviour of the source (Sect.
+4), and model its broadband emission within AGN radiative models (Sect.
 5).
-The main results are discussed and summarised in Sects.
+We discuss and summarise the main results in Sects.
 6 and 7.
 2 Observations and data reduction
-2.1 VHE -ray data: LST-1
-LST-1 is the first of the four Large-Sized Telescopes that will be constructed in the northern site of the Cherenkov Telescope Array Observatory (CTAO, Acharyya et al., 2019) at the Roque de los Muchachos Observatory in the Canary Island of La Palma (Spain).
-LST-1 is currently performing commissioning and scientific observations, after the completion of its construction in October 2018.
-Its large collection area (23 m diameter mirror) and very sensitive camera — composed of 1855 high quantum efficiency photomultiplier tubes — allow to detect faint Cherenkov light flashes from -ray showers down to energies 20 GeV (see Abe et al., 2023).
-LST-1 observed BL Lac for 23 nights between September 21 and November 25, 2022 in wobble observing mode (Fomin et al., 1994), providing 36 hours of data after applying standard data quality cuts in a zenith range between 10∘ and 60∘.
-The analysis of the raw LST-1 data was performed with version v0.10.18 of the cta-lstchain software (López-Coto et al., 2022; Moralejo et al., 2025), including data cleaning, image parametrisation, -hadron separation and direction and energy reconstruction of the recorded events.
-Among this dataset, 29 hours of data were taken under dark sky conditions (no increased background light contamination due to the presence of the moon) with a standard tail-cut cleaning with thresholds of 8 and 4 photoelectrons (p.e.) for the core and neighbour pixels of the Cherenkov shower image and an intensity cut of 50 p.e.
+2.1 Very-high-energy -ray data: LST-1
+The LST-1 is the first of the four Large-Sized Telescopes to be constructed at the northern site of the Cherenkov Telescope Array Observatory (CTAO; Acharyya et al., 2019), located at the Roque de los Muchachos Observatory on the Canary Island of La Palma (Spain).
+The LST-1 is currently performing commissioning and scientific observations, following the completion of its construction in October 2018.
+Its large collection area (23 m diameter mirror) and highly sensitive camera, composed of 1855 high quantum efficiency photomultiplier tubes, allows detection of faint Cherenkov light flashes from -ray showers down to energies 20 GeV (see Abe et al., 2023).
+The LST-1 observed BL Lac for 23 nights between 21 September and 25 November 2022 in wobble observing mode (Fomin et al., 1994), providing 36 hours of data after standard data quality cuts in a zenith range between 10∘ and 60∘.
+We analysed raw LST-1 data with version v0.10.18 of the cta-lstchain software (López-Coto and others, 2022; Moralejo et al., 2025), including data cleaning, image parametrisation, -hadron separation, and reconstruction of the direction and energy of the recorded events.
+Among this dataset, 29 hours were taken under dark-sky conditions (no increased background light contamination from the Moon), using standard tail-cut cleaning with thresholds of 8 and 4 photoelectrons (p.e.) for the core and neighbour pixels of the Cherenkov shower image and an intensity cut of 50 p.e.
 (Abe et al., 2023).
-The remaining 7 hours of observation were performed with different moon brightness levels.
+The remaining 7 hours of observation covered different moon brightness levels.
 These data were analysed with cleaning thresholds ranging from 10 to 16 p.e.
-and 5 to 8 p.e.
-for the core and neighbour pixels, and increased intensity cuts between 63 and 194 p.e.
-due to the noise introduced by the moon, extracted from the data at each night sky background (NSB) level.
-The event classification was performed with a Random Forest trained using Monte Carlo (MC) simulated proton- and -like events at a declination of 34.74∘, close to that of BL Lac (42.28∘).
-The MCs were produced to match the NSB level of our data, accounting for the additional noise with respect to dark observations.
-For the event selection, energy-dependent gammaness (gamma-likeness score) and (angular distance of the detected event with respect to the nominal source position) cuts were used, optimised to preserve 70% of the simulated MC -ray events.
-The high level analysis was done with version 1.2 of Gammapy (Donath et al., 2023; Aguasca-Cabot et al., 2023; Acero et al., 2024).
-The background was characterised using an OFF region with a wobble angle of 0.4∘ and at a reflected position of 180∘ from that of BL Lac.
-A limit on the effective area ¿5% of the maximum area was also applied.
-The statistical significance was calculated following Eq.
-(17) from Li & Ma (1983).
-The energy threshold of the light curve, calculated from the maximum of the distribution of expected events from the MCs, weighted with BL Lac’s spectrum, with the highest NSB level and zenith distance between 40∘ and 50∘ to match that of the data, was estimated to be 175 GeV.
-Therefore, we have used a minimum energy of 200 GeV for the light curve reconstruction.
-The energy threshold was further verified on data against background asymmetries in the field of view (FoV).
-We observed a non-uniformity of the event distribution, in particular along the altitude axis.
-This effect increases when the telescope pointing is further from the zenith, and becomes clearly visible at zenith distances ¿40∘.
-Looking at the energy distribution of events in the source region and the OFF region, a clear difference in energy threshold is visible for some observations.
-While the difference seems mostly at energies below our analysis threshold, we also perform our spectral analysis using backgrounds counts estimation using the package BAccMod111https://github.com/mdebony/BAccMod.
-This package allows to extract the background distribution using the full FoV after removing the source region and can thus account for asymmetries.
-As expected, we observe differences in the extracted spectra below 200 GeV for some time periods, but the spectra agree above 200 GeV thus validating our results above this energy.
-The VHE -ray light curve is shown in the top panel of Fig.
+for core pixels and 5 to 8 p.e.
+for neighbour pixels, and with increased intensity cuts between 63 and 194 p.e.
+due to noise introduced by the Moon, extracted for each night-sky background (NSB) level.
+The event classification used a Random Forest trained using Monte Carlo (MC) simulations of proton- and -like events at a declination of 34.74∘, close to that of BL Lac (42.28∘).
+We produced MC simulations to match the NSB level of our data, accounting for additional noise relative to dark observations.
+For event selection, we applied energy-dependent gammaness (gamma-likeness score) and (angular distance of the detected event with respect to the nominal source position) cuts, optimised to retain 70% of the simulated MC -ray events.
+We performed the high-level analysis with version 1.2 of Gammapy (Donath et al., 2023; Aguasca-Cabot et al., 2023; Acero et al., 2024).
+We characterised the background using an OFF region with a wobble angle of 0.4∘ and at a reflected position of 180∘ from BL Lac.
+We also applied a limit on the effective area ¿5% of the maximum area.
+We calculated the statistical significance following Eq.
+(17) from Li and Ma (1983).
+We estimated the light-curve energy threshold to be 175 GeV, calculated from the maximum of the expected event distribution from the MC simulations and weighted with BL Lac’s spectrum for the highest NSB level and zenith distance of ∘ to match the data.
+Therefore, we used a minimum energy of 200 GeV for light curve reconstruction.
+We further verified the energy threshold on the data against background asymmetries in the field of view (FoV).
+We observed a non-uniformity in the event distribution, particularly along the altitude axis.
+This effect increases as the telescope pointing is further from the zenith and becomes clearly visible at zenith distances ¿40∘.
+Examining the energy distribution of events in the source region and the OFF regions, we observe clear differences in the energy threshold for some observations.
+Although these differences occur mainly at energies below our analysis threshold, we also performed the spectral analysis using background counts estimated with the BAccMod package111https://github.com/mdebony/BAccMod.
+This package extracts the background distribution using the full FoV after removing the source region and can thus account for asymmetries.
+As expected, we observe differences in the extracted spectra below 200 GeV for some time periods, but the spectra agree above 200 GeV, validating our results at higher energies.
+We show the VHE -ray light curve in the top panel of Fig.
 1.
-We highlight October 20 and November 13 as two remarkably bright nights, when significances of 49 and 30 were reached in only 2.5 and 1.5 hours of observation.
-The minute-scale variability of these two nights is evaluated in Sect.
+We highlight October 20 and November 13, 2022 as two remarkably bright nights, when significances of 49 and 30 were reached in only 2.5 and 1.5 hours of observation, respectively.
+We evaluated the minute-scale variability of these two nights in Sect.
 3.
-In particular for October 20, consisting only of dark-sky data, we adopted a threshold of 100 GeV obtained from the dark-sky MCs with zenith 20-40∘.
-The spectrum and SED were characterised with a log-parabola model plus a term accounting for the Extragalactic Background Light (EBL) absorption with the model from Saldana-Lopez et al.
+For October 20, consisting only of dark-sky data, we adopted a threshold of 100 GeV, obtained from the dark-sky MC simulations at zenith angles 20-40∘.
+We characterised the spectrum and SED with a log-parabola model, including a term for extragalactic background light (EBL) absorption using the model from Saldana-Lopez et al.
 (2021).
-The spectral analysis is presented in detail in Sect.
+We present the spectral analysis in detail in Sect.
 4.
-2.2 HE -ray data: Fermi-LAT
-We also analysed high-energy (HE, MeV) data taken by the pair-conversion Large Area Telescope (LAT) on board the Fermi satellite.
-The LAT operates in sky-survey mode in an energy range between 20 MeV and 1 TeV, with its best sensitivity around a few GeV (Atwood et al., 2009).
-We considered a 20∘-radius region of interest (ROI) centred at the position of BL Lac.
-All data from September 21 to November 30 were downloaded, selecting the Pass8 P8R3_SOURCE_V3 events between 100 MeV and 300 GeV with the Fermitools standard software.
-A zenith angle cut of 90∘ was applied to minimise the contamination from the limb of the Earth.
-Furthermore, we used the recommended versions of the Galactic diffuse emission model and isotropic component for the event selection considered222https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html.
-To account for the emission of sources nearby BL Lac, we derived a model including all sources within the ROI, plus all sources located in an additional annular region of radius 10∘.
-This model was constructed using 1 year of data and a binned likelihood analysis procedure, using as reference the 4FGL-DR2 catalogue (Abdollahi et al., 2020; Ballet et al., 2020).
-The spectral parameters (index and normalisation) of all variable sources — as defined in the 4FGL catalogue based on the LAT variability index — contained in the ROI and with a Test Statistics (TS, Mattox et al., 1996) ¿25 (5 significance) were left as free parameters, while those not fulfilling these conditions were fixed to the catalogue values.
-The normalisation of the diffuse components was also left free.
-Convergence to the final model was achieved after removing all sources with TS¡4 (¡2).
-The HE -ray light curve was then obtained analysing the data within the desired time window through an unbinned likelihood analysis of events contained in 12-hour bins.
-For each bin we have assumed a power law spectral shape for BL Lac, leaving the spectral index and the normalisation free.
-The spectral parameters of the rest of the sources were fixed to the values from the model, with the exception of two highly variable and bright sources with a significant detection (TS¿25) located at a distance ¡15∘ from BL Lac (4FGL J2244.2+4057 and 4FGL J2311.0+3425).
-The HE -ray light curve is shown in Fig.
-1.
+2.2 High-energy -ray data: Fermi-LAT
+We analysed high-energy (HE, MeV) data taken by the pair-conversion Large Area Telescope (LAT) on board the Fermi satellite.
+The LAT operates in sky-survey mode over an energy range of 20 MeV and 1 TeV, with its best sensitivity around a few GeV (Atwood et al., 2009).
+We defined a 20∘-radius region of interest (ROI) centred on BL Lac.
+We downloaded all data from 21 September to 30 November, selecting the Pass8 P8R3_SOURCE_V3 events between 100 MeV and 300 GeV using the Fermitools standard software.
+We applied a zenith angle cut of 90∘ to minimise contamination from the Earth’s limb.
+We also applied the recommended versions of the Galactic diffuse emission model and the isotropic component for event selection222https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html.
+To account for emission from sources near BL Lac, we derived a model that includes all sources within the ROI, plus all sources located in an additional annular region of radius 10∘.
+We constructed this model using one year of data and a binned likelihood analysis, using the second release of the Fermi-LAT Fourth Source Catalogue (4FGL-DR2) as a reference (Abdollahi et al., 2020; Ballet et al., 2020).
+We left the spectral parameters (index and normalisation) of all variable sources within the ROI and with a test statistics (TS; Mattox et al., 1996) ¿25 (5 significance) as free parameters, based on the 4FGL catalogue of LAT variability.
+We fixed the sources’ parameters not fulfilling these conditions to the catalogue values.
+We also left the normalisation of the diffuse components free.
+The final model converged after removing all sources with TS¡4 (¡2).
+We obtained the HE -ray light curve by analysing the data within the chosen time window using an unbinned likelihood analysis of events in 12-hour bins.
+For each bin, we assumed a power-law spectral shape for BL Lac, leaving the spectral index and normalisation free.
+We fixed the spectral parameters of the other sources to the model value, except for two highly variable and bright sources with significant detections (TS¿25) located within ¡15∘ of BL Lac (4FGL J2244.2+4057 and 4FGL J2311.0+3425).
+Fig.
+1 shows the HE -ray light curve.
 2.3 X-ray data: Swift-XRT
-The scientific payload onboard the Neil Gehrels Swift observatory includes by the X-Ray Telescope (XRT, Burrows et al., 2005), and the Optical/UV Telescope (UVOT, Roming et al., 2005).
+The scientific payload on board the Neil Gehrels Swift observatory includes the X-Ray Telescope (XRT; Burrows and others, 2005), and the Optical/UV Telescope (UVOT; Roming and others, 2005).
 In standard operation mode, these instruments perform simultaneous observations on the same targets.
-In this project, we included 50 Swift-XRT and Swift-UVOT observations on BL Lac carried out since March 2022 up to January 2023.
+In this project, we included 50 Swift-XRT and Swift-UVOT observations of BL Lac, carried out between March 2022 and January 2023.
 Swift-XRT is sensitive to X-rays in the energy range.
-The reduced XRT data were downloaded from the UK Swift Science Data Center (Evans et al., 2009).
-The spectral analysis on the reduced data was performed using the XSPEC v12.13.1 package through PyXspec v2.1.2.
-We analysed the XRT data in the energy range from 0.3 to 10 keV.
-The intrinsic source emission in this range was primarily modelled with a power-law function and, to test the presence of intrinsic spectral curvature related to the transition between the synchrotron and IC bumps, with a log-parabola model.
-The observed spectra were fitted with the intrinsic emission models folded with the tbabs333https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node273.html Galactic absorption term, as implemented in XSPEC.
+We downloaded the reduced XRT data from the UK Swift Science Data Centre (Evans et al., 2009).
+We performed the spectral analysis of the reduced data using XSPEC v12.13.1 package through PyXspec v2.1.2.
+We analysed the XRT data in the 0.3 to 10 keV energy range.
+We primarily modelled the intrinsic source emission in this range with a power-law function, and to test for intrinsic spectral curvature associated with the transition between the synchrotron and IC bumps, we also employed a log-parabola model.
+We fitted the observed spectra with the intrinsic emission models folded with the tbabs333https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node273.html Galactic absorption term, as implemented in XSPEC.
 Following Raiteri et al.
-(2009), in the fit procedures we fixed the neutral hydrogen column density toward BL Lac to and we employed this value to correct the X-ray fluxes and SEDs for the Galactic extinction effects.
-The test was employed to assess whether the log-parabola model was preferred over the power-law one, with a significance level cut.
-The results of the X-ray analysis are summarised in Appendix B.
+(2009), we fixed the neutral hydrogen column density toward BL Lac at and used this value to correct the X-ray fluxes and SEDs for Galactic extinction effects.
+We employed the test to assess whether the log-parabola model is preferred over the power-law model, using a significance level threshold.
+We summarise the results of the X-ray analysis in Appendix B.
 2.4 Ultraviolet data: Swift-UVOT
-The analysis of UVOT data was carried out using the uvotimsum and uvotsource tasks within the HEAsoft 6.32 package, along with the latest (v20240201) Swift/UVOT Calibration Database (CALDB) files.
-UVOT performed photometric observations on BL Lac using the optical vv, bb, uu and UV w1, m2, w2 filters (Poole et al., 2008).
-Through uvotsource, we estimated the observed fluxes in the UVOT filters using a circular source region with radius of 5 centred on BL Lac and a circular background region with a 10 radius in a nearby region devoid of sources.
-The observed fluxes were corrected for Galactic extinction effects using (Schlafly & Finkbeiner, 2011) using the standard procedure described in Fitzpatrick (1999).
-Moreover, the contributions of the host galaxy in the UVOT photometric filters vv, bb, uu, w1, m2, w2, given the adopted aperture, were estimated as 2.89, 1.30, 0.36, 0.026, 0.020, and 0.017 mJy, respectively (Raiteri et al., 2013).
-Both host galaxy contributions and Galactic extinction effects were accounted for to reconstruct the intrinsic BL Lac spectral points in the optical-UV band.
-The results of the analysis of the Swift-UVOT data from the 2022 high state are reported in Table 7.
+We analysed the UVOT data using the uvotimsum and uvotsource tasks within the HEAsoft 6.32 package, along with the latest (v20240201) Swift/UVOT calibration database (CALDB) files.
+The UVOT obtained photometric observations on BL Lac through the optical vv, bb, uu and UV w1, m2, w2 filters (Poole et al., 2008).
+We estimated the observed fluxes in the UVOT filters using uvotsource, adopting a circular source region with a radius of 5 centred on BL Lac and a circular background region with a 10 radius in a nearby source-free region.
+We corrected the observed fluxes for Galactic extinction effects using (Schlafly and Finkbeiner, 2011), following the standard procedure described in Fitzpatrick (1999).
+We estimated the host galaxy contributions in the UVOT photometric filters vv, bb, uu, w1, m2, and w2, given the adopted aperture, to be 2.89, 1.30, 0.36, 0.026, 0.020, and 0.017 mJy, respectively (Raiteri et al., 2013).
+We accounted for both the host galaxy contribution and Galactic extinction effects to reconstruct the intrinsic BL Lac spectral points in the optical-UV band.
+Table 7 lists the Swift-UVOT data from the 2022 high state.
 2.5 Optical data
-Optical observations in the Johnson-Cousins bands were taken by the Sierra Nevada Observatory (SNO) as part of the TOP-MAPCAT photo-polarimetric blazar monitoring programme (Agudo et al., 2012).
-The data were reduced with the IOP4 software (Escudero Pedrosa et al., 2024a, b).
-Observations in the band were also performed as part of the Tuorla blazar monitoring444https://tuorlablazar.utu.fi/ with the 80 cm Joan Oró Telescope at Montsec Astronomical Observatory (Spain), reduced as detailed in Nilsson et al.
+We obtained optical observations in the Johnson-Cousins bands at Sierra Nevada Observatory (SNO) as part of the TOP-MAPCAT photo-polarimetric blazar monitoring programme (Agudo et al., 2012).
+We reduced the data using the IOP4 software (Escudero Pedrosa et al., 2024a, b).
+We also obtained observations in the band as part of the Tuorla blazar monitoring programme444https://tuorlablazar.utu.fi/ using the 80 cm Joan Oró Telescope at Montsec Astronomical Observatory (Spain), and reduced the data as detailed in Nilsson et al.
 (2018).
-Data in the bands were taken by the IAC80 telescope at the Teide Observatory.
-Finally, observations with the robotic telescope network of Las Cumbres Observatory (Brown et al., 2013) were also coordinated in the optical bands.
-Additionally, data from the public all-sky survey programmes and databases All-Sky Automated Survey for Supernovae (ASAS-SN, Shappee et al., 2014; Kochanek et al., 2017) in the Sloan band, and Zwicky Transient Facility (ZTF, Bellm et al., 2019) in the Sloan bands, were also included.
-These Sloan magnitudes were converted into the Johnson-Cousins photometric system using the transformations from Lupton (2005)555https://classic.sdss.org/dr4/algorithms/sdssUBVRITransform.php#Lupton2005.
-The contribution of the host galaxy emission to the total observed optical emission contained in the employed aperture has been subtracted (see Raiteri et al., 2009).
-Moreover, all data have been corrected from the reddening effect introduced by the Galactic extinction.
-For this, we adopt the extinction values , , and reported by Schlafly & Finkbeiner (2011).
-The optical flux density light curves are presented in Fig.
-1.
+Additional data were obtained in the bands with the IAC80 telescope at the Teide Observatory.
+We coordinated observations in the optical bands with the Las Cumbres Observatory robotic telescope network (Brown et al., 2013).
+We also included data from the public all-sky survey programmes and databases: the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al., 2014; Kochanek et al., 2017) in the Sloan band, and the Zwicky Transient Facility (ZTF; Bellm et al., 2019) in the Sloan bands.
+We converted these Sloan magnitudes into the Johnson-Cousins photometric system using the transformations from Lupton (2005)555https://classic.sdss.org/dr4/algorithms/sdssUBVRITransform.php#Lupton2005.
+We subtracted host galaxy contributions to the observed optical emission within our aperture (see Raiteri et al., 2009), and we corrected all data for Galactic extinction reddening.
+For this, we adopted the extinction values , , , and reported by Schlafly and Finkbeiner (2011).
+Fig.
+1 presents the optical flux density light curves.
 2.6 Radio-millimetric data
-Observations in the 37-GHz radio band were performed with the 13.7-m diameter Metsähovi radio telescope as part of its long-term AGN monitoring666https://www.aalto.fi/en/metsahovi-radio-observatory/active-galaxies (Teraesranta et al., 1998).
-Additionally, we compiled five observations in the 225-GHz band performed with the Submillimeter Array (SMA, Ho et al., 2004) at Mauna Kea, Hawaii.
-SMA is an 8-element 6-meter diameter dish radio interferometer with two orthogonally polarised receivers, allowing us to measure the total and the polarised flux of astrophysical objects.
-These receivers are inherently linearly polarised, however they can be converted to circular thanks to quarter-wave plates installed on the polarimeter (Marrone & Rao, 2008).
-These observations were carried out by the SMA Monitoring of AGNs with Polarisation (SMAPOL) programme (Myserlis et al., 2025), dedicated to the regular monitoring of -ray loud blazars, including BL Lac, on a bi-weekly cadence.
-The data were calibrated with the MIR777https://lweb.cfa.harvard.edu/~cqi/mircook.html package following standard procedures.
+The 13.7-m diameter Metsähovi radio telescope obtained 37-GHz radio band observations as part of its long-term AGN monitoring programme 666https://www.aalto.fi/en/metsahovi-radio-observatory/active-galaxies (Teraesranta et al., 1998).
+Additionally, we compiled five observations in the 225-GHz band, obtained with the Submillimeter Array (SMA; Ho et al., 2004) at Mauna Kea, Hawaii.
+The SMA is an eight-element 6-metre-diameter dish radio interferometer with two orthogonally polarised receivers, which enables the measurement of both the total and polarised flux of astrophysical objects.
+These receivers are inherently linearly polarised; however, they can be converted to circular using quarter-wave plates installed on the polarimeter (Marrone and Rao, 2008).
+These observations came from the SMA Monitoring of AGNs with Polarisation (SMAPOL) programme (Myserlis et al., 2025), which monitors -ray loud blazars, including BL Lac, on a bi-weekly cadence.
+We calibrated the data using the MIR777https://lweb.cfa.harvard.edu/~cqi/mircook.html package following standard procedures.
 3 VHE intranight variability
-In addition to the day-to-week variability, we have evaluated the VHE -ray intranight variability, a feature less often observed in blazars (other examples are e.g.
+In addition to day-to-week variability, we evaluated the VHE -ray intranight variability, a feature less often observed in blazars (other examples are e.g.
 Aharonian et al., 2007; MAGIC Collaboration et al., 2020).
-Its detection presents significant constrains to the emitting region size and the interpretation of the emission, potentially favouring leptonic models as hadrons usually require longer times for being accelerated and have longer cooling times by a factor .
-We evaluate the nights with the highest VHE -ray flux, October 20 and November 13.
-The 5-minute binned light curves are shown in Fig.
-2.
-3.1 October 20th flare
-We have evaluated the fastest variability timescales observed during the night of October 20.
-For this night we consider the integrated flux above an energy of 100 GeV, as all the data were taken under dark-sky conditions and zenith distances 20-40∘.
-As shown in the left panel of Fig.
-2, the rise and decay of the flare is well sampled by the LST-1 data, varying from 1.0 C.U.
+Its detection places significant constrains on the size of the emitting region and on the interpretation of the emission, potentially favouring leptonic models, as hadrons usually require longer acceleration times and have longer cooling times by a factor of .
+We evaluated the nights with the highest VHE -ray flux: October 20 and November 13.
+Fig.
+2 shows the 5-minute binned light curves.
+3.1 October 20 flare
+We evaluated the fastest variability timescales observed during the night of October 20.
+For this night, we considered the integrated flux above 100 GeV, as all data were taken under dark-sky conditions and zenith distances of 20-40∘.
+The left panel of Fig.
+2 shows that the LST-1 data well sample the rise and decay of the flare, varying from 1.0 C.U.
 to 4.4 C.U.
-in hours.
-This well defined rise and decay profile of the flare during this night enabled us to model the light curve as a rising and decaying exponential function, commonly used for modelling light curves (Norris et al., 1996) including those from blazar flares (see e.g.
-Aharonian et al., 2007),
-where represents the flux of the flare at a time , is the baseline flux that we have fixed at cm-2 s-1, obtained from the stable bins at the beginning and end of the flare, and and are the timescales of the flare rise and decay variations by a factor of .
-We have modelled the 5-minute binned light curve of October 20 as a sum of three possible substructures during the development of the flare with peak values found at min, min and min, considering the first flux point as min.
-We note that we have also considered the simpler cases of one or two components, comparing the goodness of the fit in each case.
-We measure values for the triple-, double- and single-exponential profiles of , and , which yields a significance of the triple-exponential model over the double- and single-function of and , respectively, using Wilks theorem (Wilks, 1938).
-We also checked that the light curve modelling does not improve adding a fourth component, with a preference over the three-component one of 0.2.
+over hours.
+This well-defined rise and decay profile of the flare allowed us to model the light curve with a rising and decaying exponential function, commonly used for modelling light curves (Norris et al., 1996), including those from blazar flares (see e.g.
+Aharonian et al., 2007), as
+Here, represents the flux of the flare at time , is the baseline flux fixed at cm-2 s-1 from the stable bins at the start and end of the flare, and and are the timescales of the flare rise and decay variations by a factor of .
+We modelled the 5-minute binned light curve of October 20 as a sum of three possible substructures during the development of the flare, with peak times at min, min, and min, considering the first flux point as min.
+We also considered the simpler cases of one or two componentsand compared the goodness of fit in each case.
+We measured values for the triple-, double-, and single-exponential profiles of , , and , which yields significances of and for the triple-exponential model over the double- and single-exponential models, respectively, using Wilks’ theorem (Wilks, 1938).
+We also verified that adding a fourth component does not improve the fit, showing a preference of only 0.2 over the three-component model.
 Therefore, we adopted the model based on the sum of three exponential profiles defined by Eq.
 (1).
-The fitted function is shown in the left panel of Fig.
-2, and the calculated rise and decay timescales of exponential profile are reported in Table 1.
-These timescales can be used to estimate the doubling flux variability timescale during the rise and decay, representative of the fastest significant variability timescales, as
-With this approach we are able to resolve variability timescales of 18 min with a confidence level of ¿, and faster hints of between 2.2 min and 4.6 min, however with a significance 2.
+We show the fitted function in the left panel of Fig.
+2, and we report the calculated rise and decay timescales of the exponential profile in Table 1.
+These timescales can be used to estimate the doubling flux variability timescale during the rise and decay, which represents the fastest significant variability timescales, as
+With this approach, we resolved variability timescales of 18 min with a confidence level of ¿, and detected faster hints between 2.2 min and 4.6 min, albeit with a significance 2.
 | Date | ||||||||
@@
 | November 13 | 0 | – | – | – |
-Alternatively, we can also estimate the fastest variability timescales through the doubling time criterion considering flux variations between consecutive temporal bins, following the procedure from Zhang et al.
-(1999) as
+We also estimated the fastest variability timescales using the doubling-time criterion for flux variations between consecutive temporal bins, following Zhang et al. (1999):
 where and are the fluxes at times and .
-We consider only those timescales obtained through this method for which the significance of the variability is 2.
-The results derived with this approach are reported in the last column of Table 1.
-We observe significant doubling flux timescales of min.
-Applying causality arguments, it is expected that very fast variability originates in very compact jet regions.
-Therefore, the fastest variability timescales observed can be used to constrain the size of the emitting region responsible for such variations as
-where is the radius of the emitting region, represents the fastest variability timescale observed, is the speed of light, is the Doppler factor of the region and corresponds to the redshift of the source (see Dondi & Ghisellini, 1995; Tavecchio et al., 2010).
-Here we adopt a conservative approach, using the shortest timescale with a significance ¿ before trials, this is, min.
-With this timescale, and considering a range of Doppler factor values , the size of the emitting region responsible for the fast variability must be cm to cm.
-We note that if the 2.2- to 6.7-minute timescales are confirmed, this would imply a region a factor 1.5 to 4 more compact.
-Nevertheless, in addition to the low significance, when considering the typical systematic errors for LST-1 flux measurements (see Abe et al., 2023), these timescales get hindered within uncertainties.
-3.2 November 13th flare
-The same analysis was performed for the night of November 13, when we also observed intranight variability in the VHE -ray emission of BL Lac.
-In this case, and owing to the higher energy threshold introduced by the moon conditions, we consider the flux above 200 GeV.
+We considered only timescales with variability significance 2.
+We list the results derived with this approach in the last column of Table 1.
+This yields a significant doubling flux timescale of min.
+Causality arguments imply that very fast variability originates in compact jet regions.
+The observed fastest variability timescales thus constrain the size of the emitting region responsible for these variations as
+where is the radius of the emitting region, represents the fastest variability timescale observed, is the speed of light, is the Doppler factor of the region, and corresponds to the redshift of the source (see Dondi and Ghisellini, 1995; Tavecchio et al., 2010).
+We adopted a conservative approach, using the shortest timescale with a significance ¿ before trials, namely min.
+With this timescale and considering a range of Doppler factor values , the size of the emitting region responsible for the fast variability is constrained to cm to cm.
+The min timescales, if confirmed, imply a region a factor 1.5 to 4 times more compact.
+Nevertheless, due to their low significance and the typical systematic errors for LST-1 flux measurements (see Abe et al., 2023), these timescales remain unconfirmed.
+3.2 November 13 flare
+We performed the same analysis for the night of November 13, when we also observed intranight variability in the VHE -ray emission of BL Lac.
+In this case, owing to the higher energy threshold introduced by moonlight, we considered the flux above 200 GeV.
 During this night, the emission decayed from 2.8 C.U.
 to 1.0 C.U.
-in 1.5 hours.
-The corresponding 5-minute binned light curve is shown in the right panel of Fig.
+over 1.5 hours.
+We show the corresponding 5-minute binned light curve in the right panel of Fig.
 2.
-In this case, LST-1 covered only the decay of the VHE -ray flare, that we modelled with a decaying exponential function as
-that follows the same description as Eq.
+The LST-1 covered only the decay of the VHE -ray flare, which we modelled with a decaying exponential function as
+following the form of Eq.
 (1).
-A value of cm-2 s-1 was used for , where the VHE flux seems to stabilise after the flare.
-The results for this night are reported as well in Table 1.
-With the estimated timescale min, we have again constrained the size of the emitting region with Eq.
-(4), for typical Doppler factors .
-This leads to a size constraint ranging from cm to cm.
-We note that in this case, the doubling flux criterion does not find significant minute-scale variability.
+We used cm-2 s-1, where the VHE flux appeared to stabilise after the flare.
+We also report the results for this night in Table 1.
+Using the estimated timescale min, we again constrained the size of the emitting region with Eq.
+(4) for typical Doppler factors .
+This constrains the size between cm and cm.
+In this case, the doubling flux criterion yields no significant minute-scale variability.
 4 Characterisation of the HE-VHE -ray spectrum
@@
 | BB10 | 59908.5 – 59909.5 | 2.3 |
-In order to evaluate possible -ray spectral variability patterns, we have studied the evolution of the -ray spectrum during the flare.
-For this, we have applied a Bayesian Block (BB) analysis on the VHE -ray light curve using the algorithm developed by Wagner et al.
-(2022)888https://github.com/swagner-astro/lightcurves, identifying 10 BBs of comparable -ray activity with a false positive rate of 1% (see Fig.
+To evaluate possible -ray spectral variability patterns, we studied the evolution of the -ray spectrum during the flare.
+For this, we applied a Bayesian block (BB) analysis on the VHE -ray light curve using the algorithm developed by Wagner et al.
+(2022)888https://github.com/swagner-astro/lightcurves, identifying ten BBs of comparable -ray activity with a false-positive rate of 1% (see Fig.
 1).
-Then, making use of the LAT and LST-1 data, we characterise the high-energy SED peak for the different identified periods from 100 MeV up to a few TeV.
-For this, we perform a joint fit of the data from both instruments with Gammapy, that allows to combine multi-instrument data at an event level, correctly accounting for instrument response functions (IRFs), backgrounds and other effects such as EBL absorption (Nievas Rosillo et al., 2025).
-For the Fermi-LAT data of each BB, we filter out all sources in the model that have a TS¡10.
-The parameters of all sources contained in a circular region of 10∘ radius centred at the position of BL Lac and with a TS¿30 are left free, while for those sources within 5∘ from BL Lac with TS¡30, only the spectral normalisation is left as a free parameter.
-For the remaining sources, all parameters are fixed to the catalogue values.
-Then, a pre-fit of the Fermi-LAT data is performed, allowing to remove weak sources still included in the model for which the number of expected counts is ¡5.
-Finally, we combine the LAT and LST-1 datasets of each BB and calculate the complete -ray spectrum and SED peak, incorporating into the analysis the response of each instrument.
-We have tested several spectral shapes of the HE-VHE emission, including log-parabola, power-law with exponential cut-off and log-parabola with exponential cut-off spectral models.
+Then, using the LAT and LST-1 data, we characterised the high-energy SED peak for the different identified periods from 100 MeV up to a few TeV.
+We performed a joint fit of the data from both instruments with Gammapy, combining multi-instrument data at the event level, while accounting for instrument response functions (IRFs), backgrounds, and other effects such as EBL absorption (Nievas Rosillo et al., 2025).
+For the Fermi-LAT data for each BB, we filtered out all sources in the model with TS¡10.
+We left free the parameters of all sources within a circular region of 10∘ radius centred at the position of BL Lac and with TS¿30, while for sources within 5∘ of BL Lac with TS¡30, we left only the spectral normalisation as a free parameter.
+For the remaining sources, we fixed all parameters to the catalogue values.
+We then performed a pre-fit of the Fermi-LAT data to remove weak sources still included in the model for which the expected counts were ¡5.
+Finally, we combined the LAT and LST-1 datasets for each BB and calculated the complete -ray spectrum and SED peak, incorporating the response of each instrument into the analysis.
+We tested several spectral shapes of the HE-VHE emission, including log-parabola, power-law with exponential cut-off and log-parabola with exponential cut-off spectral models.
 We find that the first function provides the most accurate fit to our data.
-Therefore, we use a model consisting on a log-parabola function plus a term accounting for the EBL absorption at redshift with the model from Saldana-Lopez et al.
+Therefore, we used a model consisting of a log-parabola function plus a term accounting for EBL absorption at redshift , following the model from Saldana-Lopez et al.
 (2021),
-To perform a direct comparison between the different spectra, we adopted an GeV for all periods.
-The best-fit spectral parameters are reported in Table 2.
-In Fig.
-3 we show an example of the joint SED for the eighth BB, that is, the night of November 13.
-The rest of the SEDs are available in Figs.
-6 and 7.
-Alternatively, owing to the very small variability in the spectral curvature, we tested a log-parabola function fixing , taken from the average value over the entire period.
+To directly compare the different spectra, we adopted GeV for all periods.
+Table 2 lists the best-fit spectral parameters.
+Fig.
+3 shows an example of the joint SED for the eighth BB (November 13).
+Figs.
+6 and 7 show the remaining SEDs.
+Due to the minimal variability in spectral curvature, we tested a log-parabola function with fixed at , based on the average value over the entire period.
 The results were consistent within uncertainties with those reported in Table 2.
-The complete characterisation of the high-energy SED peak performed in Sect 4 allows us to search for spectral variability patterns during the evolution of the flare.
-We observe a relatively stable value of the spectral index at 1 GeV around , with almost all BBs consistent within errors at a 2 level.
-Nevertheless, spectral index variations are particularly noticeable for the nights of October 20 with , and November 13 with , meaning that during these flares, particles are being accelerated to much higher energies.
-This is also visible from the high-energy peak shift, reported in Table 2.
-The IC peak shift becomes evident for the four brightest BBs, this is, the nights of October 20 and 27, and November 13 and 25, being the only ones with GeV.
-The average value of is found below 2 GeV, and during these nights we observe shifts up to values of 5.5 GeV for October 20, 6.5 GeV for November 13, and 2.3 GeV for October 27 and November 25.
-The Spearman correlation coefficient between and , and and show a high level of correlation ( and , respectively), indicating a possible harder-when-brighter behaviour.
-However, this correlation coefficient is dominated by the two brightest nights and does not account for uncertainties.
+The complete characterisation of the high-energy SED peak performed in Sect 4 allowed us to investigate spectral variability patterns during the evolution of the flare.
+The spectral index at 1 GeV remains stable around , with nearly all BBs consistent within errors at 2.
+Nevertheless, spectral index variations are particularly noticeable on the nights of October 20 () and November 13 (), indicating that particles are accelerated to considerably higher energies during these flares.
+This is also reflected in the high-energy peak shift, as reported in Table 2.
+The IC peak shift is evident for the four brightest BBs, namely the nights of October 20 and 27, and November 13 and 25, which are the only ones with GeV.
+The average value of is below 2 GeV, while during these nights we observe shifts of up to 5.5 GeV on October 20, 6.5 GeV on November 13, and 2.3 GeV on October 27 and November 25.
+The Spearman correlation coefficients between and , and between and , indicate a strong correlation ( and , respectively), suggesting a possible harder-when-brighter behaviour.
+However, this correlation is dominated by the two brightest nights and does not account for uncertainties.
 5 Broadband SEDs
-We have modelled the evolution of the broadband emission of BL Lac during the 2022 high state.
-We have considered 10 broadband SEDs, corresponding to each of the BBs identified above.
+We modelled the evolution of the broadband emission of BL Lac during the high state of 2022.
+We considered ten broadband SEDs, corresponding to each of the BBs identified above.
 Fig.
-4 shows as an example the model for the eighth BB.
-All fitted SED models are presented in Figs.
+4 shows, as an example, the model for the eighth BB.
+We present all fitted SED models in Figs.
 8 and 9.
-We note that, due to the lack of strictly simultaneous Swift data (0.5 days with respect to the LST-1 observations) to BBs 3, 4, 6, 8 and 9, for these periods we have considered a window of 1.5 days in order to have complete multi-wavelength coverage of the SEDs.
+We note that due to the absence of strictly simultaneous Swift data (0.5 days with respect to the LST-1 observations) to BBs 3, 4, 6, 8, and 9, we adopted a window of 1.5 days for these periods to ensure complete multi-wavelength coverage of the SEDs.
 | (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) | (12) | (13) | (14) |
@@
 | Core | 0.28 | 0.20 | 0.06 | 2.3 | 3.0 | 0.20 | 0.01 | 100 | 20 | 20 | 1.45 | – |
-We have considered a leptonic-only origin for the broadband emission.
-We use version v0.4.0 of the python package agnpy (Nigro et al., 2022, 2023).
-We find that SSC models based on a single emitting region are not able to accurately describe the broadband emission of BL Lac during this period.
-One-zone models based on EC scattering rise as an alternative to explain the VHE -ray emission, and have actually been invoked in the past for BL Lac due to the presence of a small and weak broad line region (BLR), using its external low-energy photons as seeds for the -ray production (MAGIC Collaboration et al., 2019).
-However, this solution was also found to be insufficient in this case to reproduce the broadband emission and variability during the 2022 high state, especially given the tight constrains on the emitting region size set by the intranight variability.
-Therefore, we consider a SSC model based on two emitting regions (Tavecchio et al., 2011), which has been used frequently to model the multi-wavelength emission of this source (see e.g.
-MAGIC Collaboration et al., 2019; Sahakyan & Giommi, 2022).
-We denote the two regions as ”core” and ”blob”, with sizes .
-We consider the regions to be co-spatial and interacting.
-This is motivated by the high correlation of the broadband emission from the optical band up to rays (Appendix D.2).
-A scheme of this layout is illustrated in Fig.
+We considered a leptonic-only origin for the broadband emission.
+We used version v0.4.0 of the python package agnpy (Nigro et al., 2022, 2023).
+We find that SSC models based on a single emitting region are unable to accurately describe the broadband emission of BL Lac during this period.
+One-zone models based on EC scattering present an alternative to explain VHE -ray emission and have previously been invoked for BL Lac due to the presence of a small and weak broad line region (BLR), which provides external low-energy photons as seeds for -ray production (MAGIC Collaboration et al., 2019).
+However, we find that this solution is insufficient to reproduce the broadband emission and variability during the 2022 high state, particularly given the tight constraints on the size of the emitting region imposed by the intranight variability.
+We therefore considered an SSC model based on two emitting regions (Tavecchio et al., 2011), which has frequently been used to model the multi-wavelength emission of this source (see e.g.
+MAGIC Collaboration et al., 2019; Sahakyan and Giommi, 2022).
+We denote the two regions as ‘core’ and ‘blob’, with .
+We consider the regions to be co-spatial and interacting, motivated by the high correlation of the broadband emission from the optical band up to rays (Appendix D.2).
+Fig.
 9b of MAGIC Collaboration et al.
-(2019).
-Each region is defined by a magnetic field , a Lorentz factor , a Dopper factor and an electron population that we have characterised through a broken power-law distribution with respect to the electron Lorentz factor ,
+(2019) illustrates a schematic of this layout.
+Each region is characterised by a magnetic field , a Lorentz factor , a Dopper factor , and an electron population described by a broken power-law distribution with respect to the electron Lorentz factor ,
 where is the normalisation of the distribution, and are the minimum and maximum values of the Lorentz factor, and and are the spectral indices below and above the break of the distribution , respectively.
-We constrain the size of the blob, responsible for the intranight VHE -ray variability, with the estimates presented in Sect.
+We constrained the size of the blob, responsible for the intranight VHE -ray variability, using the estimates presented in Sect.
 3.
-Additionally, we have incorporated to the SED model an EC contribution from external photon fields, namely the BLR, the accretion disc and the dusty torus.
-This is because, as mentioned above, these components have been used in the past to reproduce the large width of the IC SED peak (see e.g.
+Additionally, we incorporated an EC contribution from external photon fields in the SED model, namely the BLR, the accretion disc, and the dusty torus, since these components have previously been used to reproduce the broad IC SED peak (see e.g.
 Khatoon et al., 2024).
-Following Ghisellini & Tavecchio (2009); MAGIC Collaboration et al.
-(2019), we adopt erg s-1 and cm for the BLR, which leads to erg s-1 and cm using the typical relations between BLR, disc and torus (see Kaspi et al., 2007; Cleary et al., 2007), assuming in this case a covering factor for torus of , and a BLR reprocessing factor (based on the Hα line) .
+Following Ghisellini and Tavecchio (2009); MAGIC Collaboration et al.
+(2019), we adopted erg s-1 and cm for the BLR, which implies erg s-1 and cm using the typical relations between BLR, disc, and torus (see Kaspi et al., 2007; Cleary et al., 2007).
+We assumed a torus covering factor and a BLR reprocessing factor (based on the Hα line) .
 For the dusty torus, we assumed a typical temperature of 1000 K.
-We adopted a black hole mass of (Ghisellini et al., 2010).
-Finally, we considered the core located at a distance cm from the central engine and the blob at cm.
-We also take into account the possible absorption at GeV-TeV energies and the EBL absorption at VHE.
-We also evaluated the absorption due to the external radiation fields, which was found to be negligible.
-Multi-zone models of blazar SEDs are characterised by a high degree of degeneracy, owing to the relatively large number of parameters.
-This degeneracy generally prevents the use of standard fit routines based on the minimisation of a cost function to effectively constrain the parameter space.
-In such cases, it is necessary to adopt a physically motivated scenario as guiding framework.
-In this work, we have investigated a scenario in which the evolution of the flare is driven by magnetic reconnection, which is strongly supported by the minute-scale variability detected at VHE with LST-1.
-In order to fit the SED of each BB, for each block we have used the parameters of the previous one as initial parameters, iterating until reaching a good representation of the SED.
-To reduce the number of free parameters, we have kept the magnetic fields , the emitting region sizes , the Lorentz factor and the Doppler factor fixed (with values in agreement with the limit values set by the variability analysis).
-Additionally, we minimise the changes of physical parameters in the core region between adjacent BBs to reflect the slow evolving nature of the core, unless the observed time interval between the centre of two adjacent BBs is longer than two days, the light crossing timescale of the core region in our reference frame.
-The blob, in contrast, is considered as the newly formed plasmoid from an unresolved magnetic reconnection layer in the core, whose parameters can vary considerably depending on the reconnection layer.
-Therefore, we have based our SED evolution in changes of the electron distribution of each region, that is, , , , , and .
-The parameter change was done in a way that is consistent with the flare’s evolution timescales.
-The resulting models for each BB are shown in Figs.
-8 and 9, and the parameters describing each of them are summarised in Table 3.
-The consistency of these models with a magnetic reconnection scenario is discussed in Sect.
+We adopted a black hole mass (Ghisellini et al., 2010).
+We considered the core located at a distance cm from the central engine and the blob at cm.
+We accounted for possible absorption at GeV-TeV energies and the EBL absorption at VHE.
+Our analysis shows that absorption due to external radiation fields is negligible.
+Multi-zone models of blazar SEDs are characterised by a high degree of degeneracy due to the relatively large number of parameters.
+This degeneracy generally prevents the use of standard fitting routines based on cost-function minimisation for effectively constraining the parameter space.
+In such cases, it is necessary to adopt a physically motivated scenario as a guiding framework.
+In this work, we investigated a scenario in which the evolution of the flare is driven by magnetic reconnection, which is strongly supported by the minute-scale variability detected at VHE with LST-1.
+To fit the SED of each BB, we used the parameters of the previous block as initial parameters, iterating until we achieved a satisfactory representation of the SED.
+To reduce the number of free parameters, we kept the magnetic fields , the emitting region sizes , the Lorentz factor , and the Doppler factor fixed, adopting values consistent with the limit values set by the variability analysis.
+Additionally, we minimised changes in the physical parameters in the core region between adjacent BBs to reflect the slowly evolving nature of the core, unless the time interval between the centre of two adjacent BBs exceeded two days, corresponding to the light-crossing timescale of the core region in our reference frame.
+In contrast, the blob is considered as a newly formed plasmoid originating from an unresolved magnetic reconnection layer in the core, whose parameters can vary considerably depending on the reconnection layer.
+Therefore, we based our SED evolution on changes in the electron distribution of each region, namely , , , , , and .
+The parameter changes were implemented in a manner consistent with the flare evolution timescales.
+Figs.
+8 and 9 show the resulting models for each BB, and Table 3 summarises the corresponding parameters.
+We discuss the consistency of these models with a magnetic reconnection scenario in Sect.
 6.
-The leptonic origin of the emission considered here is supported by recent results obtained by the IXPE (Imaging X-ray Polarimetry Explorer) satellite, based on the analysis of the broadband polarised emission of BL Lac (see Middei et al., 2023; Peirson et al., 2023).
-Under a leptonic framework where the X-ray spectrum is dominated by the high-energy IC scattering component, the X-ray polarisation degree is expected to be substantially lower than that of the optical band, which originates from electron synchrotron radiation (Liodakis et al., 2019; Peirson & Romani, 2019; Zhang et al., 2024).
+The leptonic origin of the emission considered here is supported by recent results obtained by the Imaging X-ray Polarimetry Explorer (IXPE) satellite, based on the analysis of the broadband polarised emission of BL Lac (see Middei et al., 2023; Peirson et al., 2023).
+Within a leptonic framework in which the X-ray spectrum is dominated by the high-energy IC scattering component, the X-ray polarisation degree is expected to be substantially lower than that of the optical band, which originates from electron synchrotron radiation (Liodakis et al., 2019; Peirson and Romani, 2019; Zhang et al., 2024).
 Middei et al.
 (2023) measured an upper limit of % for two IXPE observations of BL Lac during May and July 2022, consistent with this scenario given the optical polarisation degree.
-Moreover, under strong flares where the soft X-ray emission has significant contribution from synchrotron radiation as observed from the much flatter X-ray spectrum — which is also the present case, see for instance BB3 with respect to e.g.
-BB8 when the source was at its brightest — one expects a significantly higher polarisation degree of the soft X-ray photons than that at hard X-ray energies, and even higher than the optical polarisation degree (Liodakis et al., 2019; Peirson & Romani, 2019; Zhang et al., 2024).
-A third IXPE observation of BL Lac, performed two days after the last LST-1 observation (November 27 to 29) resulted in a measured X-ray polarisation degree of over the entire X-ray band.
-After splitting the analysis in two bins accounting for the soft and hard X-ray regimes, almost all the polarised emission was found to be in the former, with a polarisation degree of .
-The synchrotron emission was estimated to comprise 44% of the total flux at the soft X-ray band (Peirson et al., 2023).
-After accounting only for the synchrotron emission and assuming that the 56% of the flux remaining comes from unpolarised SSC, the X-ray polarisation fraction was calculated to be 49%, much higher than that in optical frequencies (13%), and well within the expectations for a SSC scenario (Peirson & Romani, 2019).
-Therefore, the leptonic SED modelling performed here is consistent with the expected nature of the broadband emission based on both the spectral and the multi-wavelength polarisation analyses.
-A last IXPE campaign on BL Lac during November 2023 resulted in an extremely high optical polarisation of %, not accompanied however with a high X-ray polarisation, leading to a 3 upper limit of 7.4%.
-This was observed during a period in which BL Lac was behaving as an LSP.
-Therefore, the X-ray emission was located in the high-energy SED component.
-These results strongly favour SSC multi-zone or energy-stratified models (see Zhang et al., 2024; Agudo et al., 2025), leaving little room for hadronic contribution, only in the framework of a hybrid model as shown by a detailed analysis of the broadband emission by Liodakis et al.
+Moreover, during strong flares in which the soft X-ray emission has a significant synchrotron radiation contribution, as indicated by a much flatter X-ray spectrum (as in the present case, e.g.
+BB3 compared to BB8 when the source was at its brightest), we expect a significantly higher polarisation degree of soft X-ray photons than at hard X-ray energies, and even higher than the optical polarisation degree (Liodakis et al., 2019; Peirson and Romani, 2019; Zhang et al., 2024).
+A third IXPE observation of BL Lac, performed two days after the last LST-1 observation (27-29 November) resulted in a measured X-ray polarisation degree of over the entire X-ray band.
+After splitting the analysis into two bins corresponding to the soft and hard X-ray regimes, almost all the polarised emission appears in the soft X-ray bin, with a polarisation degree of .
+Peirson et al.
+(2023) estimated that the synchrotron emission comprises 44% of the total flux in the soft X-ray band.
+After accounting only for synchrotron emission and assuming that the remaining 56% of the flux originates from unpolarised SSC emission, these authors calculated the X-ray polarisation fraction to be 49%, much higher than that in optical frequencies (13%), and well within the expectations for an SSC scenario (Peirson and Romani, 2019).
+Therefore, the leptonic SED modelling performed here is consistent with the expected nature of the broadband emission based on both spectral and multi-wavelength polarisation analyses.
+A final IXPE campaign on BL Lac during November 2023 resulted in an extremely high optical polarisation of %, not accompanied by high X-ray polarisation, yielding a 3 upper limit of 7.4%.
+BL Lac behaved as an LSP during this period, placing the X-ray emission in the high-energy SED component.
+These results strongly favour SSC multi-zone or energy-stratified models (see Zhang et al., 2024; Agudo et al., 2025), leaving little room for a hadronic contribution, except within the framework of a hybrid model, as shown by the detailed broadband analysis of Liodakis et al.
 (2025).
-However, as presented by these authors, hybrid models expect a hardening of the VHE spectrum that is not observed here in order to reproduce the multi-wavelength polarisation signatures.
+However, as shown by these authors, hybrid models expect a hardening of the VHE spectrum to reproduce the multi-wavelength polarisation signatures, which is not observed in the present data.
 6 Discussion
 6.1 -ray and X-ray spectral behaviour
-X-ray spectral trends were also observed in BL Lac prior to the 2022 high state (e.g.
+X-ray spectral trends also appear in BL Lac prior to the 2022 high state (e.g.
 Weaver et al., 2020).
-Through an extensive SED modelling, Sahakyan & Giommi (2022) attributed the synchrotron component emerging in X-rays during periods of high state to a second emission zone located beyond the BLR, different from the one dominating the emission in the optical-UV band.
-The emission from this second region can produce a shift of the synchrotron peak towards higher frequencies, reaching values of even 10 Hz, well within the high-synchrotron-peaked (HSP) range.
-Indeed, BL Lac is well known for transitioning between the LSP regime — mainly during low flux states — to an IBL/HBL behaviour during active phases (see e.g.
+Through extensive SED modelling, Sahakyan and Giommi (2022) attribute the synchrotron component emerging in X-rays during periods of high state to a second emission zone located beyond the BLR, distinct from the region that dominates the emission in the optical-UV band.
+Emission from this second region can shift the synchrotron peak towards higher frequencies, reaching 10 Hz, well within the high-synchrotron-peaked (HSP) range.
+BL Lac is well known for transitioning between the LSP regime, primarily during low flux states, to an IBL-HBL behaviour during active phases (see e.g.
 Ackermann et al., 2011; Nilsson et al., 2018).
-Starting on approximately 2019, the source has been going through several years of very high activity, culminating with its three brightest flares ever recorded in 2021, 2022 and 2024.
-Therefore, the X-ray spectral variability trends observed in the dataset analysed in this work can be related to the emergence of synchrotron emission from freshly accelerated particles in a secondary emission region along the jet.
-The shift of the X-ray emission from the high- to the low-energy peak of the SED can also be the reason behind the lower correlation of the hard X-ray and optical flux compared to the soft X-ray one, as reported in Appendix D.2.
-Furthermore, in our model the emission from the secondary region (blob) dominates also in the HE and VHE -ray bands.
-Therefore, the same freshly accelerated particles are also responsible for the intense flaring activity detected with Fermi-LAT and LST-1 via SSC, as well as for the observed brightening and shifting of the -ray spectrum towards higher energies (see Sect.
+Since approximately 2019, the source has undergone several years of very high activity, culminating in its three brightest flares ever recorded in 2021, 2022, and 2024.
+Therefore, the X-ray spectral variability trends observed in the dataset analysed here can be attributed to the emergence of synchrotron emission from freshly accelerated particles in a secondary emission region along the jet.
+The shift of the X-ray emission from the high- to low-energy peak of the SED may also explain the lower correlation between the hard X-ray and optical flux compared to the soft X-ray flux, as reported in Appendix D.2.
+Furthermore, in our model, the emission from the secondary region (blob) also dominates in the HE and VHE -ray bands.
+Consequently, these freshly accelerated particles are responsible for the intense flaring activity detected with Fermi-LAT and LST-1 via SSC, as well as for the observed brightening and shifting of the -ray spectrum towards higher energies (see Sect.
 4).
 6.2 Week-scale evolution of the broadband emission
-We successfully model the broadband emission and evolution of the flare as a SSC two-zone model considering as well the EC contribution.
+We successfully modelled the broadband emission and evolution of the flare as an SSC two-zone model, including the EC contribution.
 The flare begins with emission dominated by the slowly evolving core, extending up to X-ray energies.
-Plasmoids (blobs) can develop within some considerably magnetised current sheets within the core through magnetic reconnection.
-These blobs become primarily responsible for the gamma-ray emission.
-At this early stage of the flare, the EC scattering also contributes to the gamma-ray emission.
-As magnetic reconnection sets in, particles in the blob are accelerated to higher energies, increasing its contribution also in the X-ray band.
-Our spectral fitting reveals several major ejections of plasmoids/blobs in the emission region, namely, BB4, BB6, BB8, and BB10.
-During all these episodes, a freshly produced blob dominates the gamma-ray emission, and can also have strong contributions in other wavelengths depending on the blob parameters.
-All those episodes are followed by a relatively low flux episode, where the newly formed blob cools and loses its energy.
-Since each blob is produced independently, their parameters are expected to be considerably different.
-On the other hand, the core evolves slowly, thus its physical parameters do not change much between different episodes.
-This layout is in agreement with the correlation analysis performed in Appendix D.2, where the high degree of correlation between different bands can be attributed to the co-spatial nature of the broadband emission, with the exception of the lower correlation found with the radio and VHE bands.
-For the former, it is known that the inner regions of the jet are highly opaque to radio emission, that is typically assumed to come from an outer part (Max-Moerbeck et al., 2014).
-Several studies have in fact found long delays of hundreds of days between the radio and multi-wavelength emission of blazars (see e.g.
+Plasmoids (blobs) can develop within highly magnetised current sheets in the core via magnetic reconnection.
+These blobs are primarily responsible for the gamma-ray emission.
+At this early stage, EC scattering also contributes to the gamma-ray emission.
+As magnetic reconnection progresses, particles in the blob are accelerated to higher energies, increasing the blob’s contribution in the X-ray band.
+Our spectral fitting reveals several major ejections of plasmoids and/or blobs in the emission region, specifically BB4, BB6, BB8, and BB10.
+During each of these episodes, a newly formed blob dominates the gamma-ray emission and may contribute significantly at other wavelengths, depending on the blob parameters.
+Each episode is followed by a relatively low-flux period during which the newly formed blob cools and loses its energy.
+Because each blob forms independently, their parameters are expected to vary considerably, whereas the core evolves slowly, so its physical parameters remain largely unchanged between different episodes.
+This scenario is consistent with the correlation analysis presented in Appendix D.2, where the high degree of correlation between different bands is attributed to the co-spatial nature of the broadband emission, with the exception of the lower correlation observed in the radio and VHE bands.
+For the radio band, it is well established that the inner regions of the jet are highly opaque to radio emission, which typically originates from an outer part (Max-Moerbeck et al., 2014).
+Several studies report long delays of hundreds of days between the radio and multi-wavelength emission of blazars (see e.g.
 Liodakis et al., 2018).
-We cannot test this hypothesis here as we focus on a shorter period.
-Nevertheless, such behaviour could explain the lower correlation of this band.
-We note, however, that the radio band as a lower observing cadence than that in optical, that can also be biasing the measured correlation.
-Regarding the latter band, the lower correlation of the VHE emission can be explained by the shorter cooling timescales of the highest-energy electrons responsible for the TeV emission.
-However, as for the radio band correlation, the lower observing cadence of the VHE data can introduce systematic effects in the correlation analysis, diluting the correlation of the TeV emission.
-With the configuration assumed in Sect.
-5, the low-energy photons from dusty torus are the ones dominating the EC contribution.
-This EC emission has been used in the past to model the broadband emission of BL Lac (Khatoon et al., 2024).
-In the present flare, owing to the very bright state in which the source was observed during most of the period, we find that the EC emission becomes relevant only at the early stages of the outburst, when the source is faintest (see the model for BB1 in Fig.
+We could not test this hypothesis as we focused on a shorter period.
+Nevertheless, such behaviour could explain the lower correlation observed in this band.
+We note, however, that the radio band has a lower observing cadence than the optical band, which can also bias the measured correlation.
+For the VHE band, the lower correlation can be explained by the shorter cooling timescales of the highest-energy electrons responsible for the TeV emission.
+However, as with the radio band, the lower observing cadence of the VHE data may introduce systematic effects in the correlation analysis that dilute the correlation in the TeV emission.
+With the configuration adopted in Sect.
+5, the low-energy photons from dusty torus dominate the EC contribution.
+This EC emission has previously been used to model the broadband emission of BL Lac (Khatoon et al., 2024).
+In the present flare, due to the very bright state of the source during most of the period, we find that the EC emission is only relevant during the early stages of the outburst, when the source is faintest (see the model for BB1 in Fig.
 8).
-We have also considered the possible emission arising from the interplay between the core and blob regions, this is, low-energy photons from one region that can be IC-scattered by electrons of the other region.
-However, this was found to be negligible due to the compactness of the blob as a result of the size constraints derived from the VHE minute-scale variability.
-The absorption was also evaluated and taken into account, but was found to be negligible up to energies TeV, where the sources is not detected significantly and only upper limits are estimated.
+We also considered possible emission arising from the interaction between the core and blob regions, that is, low-energy photons from one region being IC-scattered by electrons in the other region.
+However, this contribution was found to be negligible because of the compactness of the blob as a result of the size constraints derived from the VHE minute-scale variability.
+We also evaluated the absorption and found it to be negligible up to TeV energies, where the source is not detected significantly and we estimate only upper limits.
 Overall, the parameters are consistent with the typical values derived from a large sample of blazars (Ghisellini et al., 2010).
-The core region is mainly responsible for the emission from radio to UV wavelengths, and contributes significantly to the hard X-ray emission during those periods in which the X-ray emission is found in the transition from the low- to the high-energy peak of the SED.
-During the early moments of the flare, when the soft X-rays are also found in the IC peak, it dominates both the soft and hard X-ray emission.
-Changes in the parameters between adjacent BBs are rather small considering the time interval between consecutive blocks is generally longer than the light crossing time of the core, reflecting its slow evolution.
-On the other hand, during those periods when BL Lac fully transitions from LSP to ISP and the X-ray emission is entirely produced by synchrotron radiation, the low energy emission of the blob is the one becoming dominant in the X-ray band.
-We note that the values of the blob are on the high end of normal values.
-Zech & Lemoine (2021) interpreted these high values in other BL Lac objects as a thermal coupling of protons and electrons in a weakly magnetised shock, where protons transfer energy to electrons in the shock layer.
-The main driver of the variability on day-to-week timescales is the evolution of the particle population of each region during the flare.
-Especially noticeable is the electron distribution behaviour of the blob during the brightest nights (BB4 and BB8), when it clearly reaches higher energies, in consistency with the IC peak shift towards higher energies reported in Sect.
+The core region primarily drives the emission from radio to UV wavelengths and contributes significantly to the hard X-ray emission during periods where the X-ray emission is transitioning from the low- to high-energy peak of the SED.
+During the early stages of the flare, when the soft X-rays also lie within the IC peak, the core dominates both the soft and hard X-ray emission.
+Changes in the parameters between adjacent BBs are relatively small, given that the time interval between consecutive blocks generally exceeds the light-crossing time of the core, reflecting its slow evolution.
+Conversely, during periods when BL Lac fully transitions from LSP to ISP and the X-ray emission is entirely produced by synchrotron radiation, the low-energy emission from the blob becomes dominant in the X-ray band.
+The values of the blob are at the high end of normal values.
+Zech and Lemoine (2021) interpret such high values in other BL Lac objects as resulting from thermal coupling between protons and electrons within a weakly magnetised shock, where protons transfer energy to electrons within the shock layer.
+The primary driver of variability on day-to-week timescales is the evolution of the particle populations within each region during the flare.
+The blob’s electron distribution behaviour of the blob is particularly noticeable during the brightest nights (BB4 and BB8), when it clearly reaches higher energies, consistent with the IC peak shift towards higher energies reported in Sect.
 4.
-Moreover, for these two nights, an injection of freshly accelerated particles was also needed to explain the remarkably high -ray emission.
-The blob is also characterised by a higher magnetisation and higher Doppler boosting than the core, with G, and G, , respectively.
-This, combined with the equipartition values of the core, close to equipartition but still 1, and its sometimes very hard spectral index, are consistent with a magnetic reconnection scenario (see e.g.
-Jormanainen et al., 2023), discussed in detail in Sect.
+Furthermore, for these two nights, an injection of freshly accelerated particles was also necessary to account for the remarkably high -ray emission.
+The blob is also characterised by higher magnetisation and Doppler boosting than the core, with G, , and G, .
+This, combined with the equipartition values of the core, which are close to equipartition but still 1, and its occasionally very hard spectral index, is consistent with a magnetic reconnection scenario (see, e.g.
+Jormanainen et al., 2023), as discussed in detail in Sect.
 6.3.
-We note that the equipartition values of the blob are far from 1, with the electron energy density dominating on the order of magnitude of .
-This is a limitation that is regularly found in these models (see Tavecchio & Ghisellini, 2016).
-This might arise from assuming a still over-simplified picture of the relativistic jets while trying to capture all the observational features, including, week, day and minute scale variability.
-Nevertheless, the models presented here are able to accurately describe the evolution of the flare while still providing a reasonable solution to observational features such as multi-wavelength correlations, week-, day- and minute-scale VHE variability.
+The equipartition values of the blob are far from 1, with the electron energy density dominating on the order of magnitude of .
+This limitation is commonly encountered in such models (see Tavecchio and Ghisellini, 2016).
+This may result from adopting a an overly simplified representation of relativistic jets while attempting to capture all observational features, including, week-, day-, and minute-scale variability.
+Nevertheless, the models presented here can accurately describe the evolution of the flare while providing a reasonable solution to observational features such as multi-wavelength correlations and week-, day-, and minute-scale VHE variability.
 6.3 Origin of the minute-scale VHE variability
-The very fast minute-scale variability demands a compact region with extreme particle acceleration (the blob in our fitting model) on top of the blazar zone (the core in our model) that governs the day-to-week-scale variability.
+The extremely fast, minute-scale variability requires a compact region with extreme particle acceleration (the blob in our fitting model) in addition to the blazar zone (the core in our model) , which governs the day-to-week-scale variability.
 Such a physical setup is often attributed to the mini-jet scenario under magnetic reconnection (Giannios et al., 2009; Petropoulou et al., 2016).
-A particularly high value of the blob Lorentz factor is also necessary to avoid the VHE -ray attenuation in the compact blob.
-The optical depth for VHE -rays is given by (Ackermann et al., 2016), where is the observed VHE photon energy, is the optical luminosity, which is the target photons for VHE -rays, and is the shortest observed variability timescale.
-We find that the Doppler factor can ensure the compact blob is optically thin to the observed VHE photons.
-This range of Doppler factor values is typically high with respect to those often observed in blazars.
-Hence, the mini-jet scenario is particularly relevant here to solve this problem and avoid the VHE -ray attenuation in the compact blob.
-Under the mini-jet scenario, the terminal bulk Lorentz factor of the plasmoid can reach approximately the magnetisation factor of the reconnection site in the co-moving frame (Sironi et al., 2016).
-As shown in Table 3, the magnetisation factor in the core, estimated by , is generally .
-The bulk Lorentz factor of the blob relative to the core can be approximated by (if they are both close to the line of sight), which is comparable to the magnetisation factor.
-Although our fitting model and the above estimates are rather simplified, they are consistent with the mini-jet in magnetic reconnection scenario (Nalewajko et al., 2011).
-A plasmoid merger in magnetic reconnection can be an alternative scenario (Zhang et al., 2022).
-Previous particle-in-cell (PIC) simulations have shown that the SSC can exhibit much higher and faster variability than the synchrotron and EC, due to the highly concentrated electrons freshly accelerated at the plasmoid merging site.
-The magnetisation factor therein can drop below 1 even though the reconnection site has overall 1 magnetisation, consistent with the fitting parameters derived for the blob (Zhang et al., 2022).
-In principle, such collisions and mergers of plasma structures can happen in magnetised turbulence as well (Zhang et al., 2023).
-However, the very high intrinsic X-ray polarisation (49%, see Peirson et al., 2023) measured just two days after the last flare disfavours a turbulent environment.
-In addition to the fast variability, the very hard particle spectrum in the core () also indicates a magnetic reconnection origin.
-Although we fit the radio to optical spectrum by the synchrotron emission from the core, usually the large-scale jet can have a considerable contribution to the radio emission.
-Therefore, the radio to optical emission intrinsic to the core in our model may require even harder particle spectrum.
-It is generally believed that only magnetic reconnection in a highly magnetised environment can result in particle spectrum harder than 2 (Guo et al., 2016, 2021).
+The blob also requires a particularly high Lorentz factor to prevent VHE -ray attenuation within the compact blob.
+The optical depth for VHE -rays is given by (Ackermann et al., 2016), where is the observed VHE photon energy, is the optical luminosity providing the target photons for VHE -rays, and is the shortest observed variability timescale.
+We find that the Doppler factor ensures that the compact blob is optically thin to the observed VHE photons.
+This range of Doppler factor values is typically higher than those commonly observed in blazars.
+Hence, the mini-jet scenario is particularly relevant here to solve this problem and prevent VHE -ray attenuation within the compact blob.
+In the mini-jet scenario, the terminal bulk Lorentz factor of the plasmoid can approximately equal the magnetisation factor of the reconnection site in the co-moving frame (Sironi et al., 2016).
+Table 3 shows that the magnetisation factor in the core, estimated by , is generally .
+The bulk Lorentz factor of the blob relative to the core can be approximated by (assuming both are close to the line of sight), which is comparable to the magnetisation factor.
+Although our fitting model and the above estimates are rather simplified, they remain consistent with the mini-jet within a magnetic reconnection scenario (Nalewajko et al., 2011).
+A plasmoid merger in magnetic reconnection presents an alternative scenario (Zhang et al., 2022).
+Previous particle-in-cell (PIC) simulations show that the SSC can exhibit much higher and faster variability than the synchrotron and EC, owing to the highly concentrated electrons freshly accelerated at the plasmoid merging site.
+The magnetisation factor at such sites can drop below one, even though the reconnection site has an overall magnetisation greater than one, consistent with the fitting parameters derived for the blob (Zhang et al., 2022).
+In principle, such collisions and mergers of plasma structures can also occur in magnetised turbulence (Zhang et al., 2023).
+However, the very high intrinsic X-ray polarisation (49%, see Peirson et al., 2023), measured just two days after the last flare, disfavours a turbulent environment.
+In addition to the rapid variability, the very hard particle spectrum in the core () also indicates a magnetic reconnection origin.
+Although we fitted the radio-to-optical spectrum by the synchrotron emission from the core, the large-scale jet can often contribute significantly to the radio emission.
+Consequently, the radio-to-optical emission intrinsic to the core in our model may require an even harder particle spectrum.
+It is generally accepted that only magnetic reconnection in a highly magnetised environment can produce a particle spectrum harder than two (Guo et al., 2016, 2021).
 Interestingly, since the correlated quasi-periodic flux and polarisation variations detected by Jorstad et al.
-(2022) suggested kink instabilities in BL Lac, and the very high optical polarisation degree found in Agudo et al.
-(2025) indicates the relaxation of a magnetic spring in the jet, with the fast variability reported in the current paper, it is possible that the strong activities of BL Lac since roughly 2020 are associated with giant eruption of magnetised plasma from the central engine, likely due to the magnetically-arrested-disc jet scenario.
-Under this picture, our two-zone model essentially consists of a segment of the kinked jet (the core) and plasmoids (the blob) produced via magnetic reconnection that is triggered by the kinked jet from time to time.
-Since they are co-spatial and physically connected, all bands above the synchrotron self absorption limit should have some correlations.
-However, the production of plasmoids and their mergers in reconnection is rather random as shown in numerical simulations (Sironi et al., 2016; Zhang et al., 2022).
-Hence the VHE emission that is fully dominated by the blob can appear less correlated to the optical emission that is fully dominated by the core.
+(2022) suggest kink instabilities in BL Lac, and the very high optical polarisation degree reported by Agudo et al.
+(2025) indicates the relaxation of a magnetic spring in the jet, combined with the fast variability observed in this work, it is plausible that the intense activity of BL Lac since approximately 2020 is associated with a giant eruption of magnetised plasma from the central engine, likely arising from a magnetically-arrested-disc jet scenario.
+Within this framework, our two-zone model essentially consists of a segment of the kinked jet (the core) and plasmoids (the blob) produced by magnetic reconnection, which is periodically triggered by the kinked jet.
+As they are co-spatial and physically connected, all bands above the synchrotron self-absorption limit should exhibit some degree of correlation.
+However, the production of plasmoids and their mergers during reconnection is somewhat random, as demonstrated in numerical simulations (Sironi et al., 2016; Zhang et al., 2022).
+Consequently, the VHE emission, which is fully dominated by the blob, may appear less correlated with the optical emission, which is fully dominated by the core.
 Moreover, the models presented here propose a simplified picture of the magnetic reconnection scenario.
-As the flare develops, plasmoids grow and can be ejected from the reconnection layer with various physical conditions as suggested by recent PIC simulations (Sironi et al., 2016; Zhang et al., 2024).
-While a fully numerical treatment of plasma dynamics in reconnection is far beyond the scope of this paper, we model these plasmoids as one blob with highly variable physical parameters as shown in many previous studies (e.g.
+As the flare develops, plasmoids grow and can be ejected from the reconnection layer under various physical conditions, as indicated by recent PIC simulations (Sironi et al., 2016; Zhang et al., 2024).
+Although a fully numerical treatment of plasma dynamics during reconnection is beyond the scope of this paper, we modelled these plasmoids as a single blob with highly variable physical parameters, as demonstrated in several previous studies (e.g.
 Christie et al., 2019).
-We note that the two-zone model has significant parameter degeneracy, thus our fitting parameters may not be unique.
-Additionally, we have calculated the blob synchrotron and SSC cooling timescales in the observer frame for the electrons responsible for the GeV-TeV -ray emission.
-The synchrotron cooling timescales were estimated as (see e.g.
+We note that the two-zone model exhibits significant parameter degeneracy; consequently, our fitting parameters may not be unique.
+Furthermore, we calculated the synchrotron and SSC cooling timescales of the blob in the observer frame for the electrons responsible for the GeV-TeV -ray emission.
+We estimated the synchrotron cooling timescales as (see e.g.
 Abe et al., 2025)
-where is the electron mass, is the Thomson cross-section, is the magnetic field of the blob and is the Lorentz factor of the electrons.
-The blob SSC cooling timescales in the observer frame were computed as , with expressed in the blob reference frame as (Moderski et al., 2005)
-where is the photon energy in the rest frame of the blob in units of the electron rest mass energy, , being the photon frequency in the blob frame, is defined by Eq.
+where is the electron mass, is the Thomson cross-section, is the magnetic field of the blob, and is the Lorentz factor of the electrons.
+We calculated the SSC cooling timescales of the blob in the observer frame as , with expressed in the blob reference frame as (Moderski et al., 2005)
+Here, denotes the photon energy in the rest frame of the blob in units of the electron rest-mass energy, , with representing the photon frequency in the blob frame.
+The function is defined by Eq.
 (C5) in Moderski et al.
-(2005), is the synchrotron photon number density in the blob rest frame, computed as in Abe et al.
+(2005), is the synchrotron photon number density in the blob rest frame, calculated as in Abe et al.
 (2025).
-The blob SSC cooling timescales were found to be shorter than the synchrotron ones, consistent with the Compton dominance of – shown by the blob SED.
-In Table 3 we report the blob SSC cooling timescales at .
-For the two nights with intranight variability, our models predict SSC cooling timescales of the order of few minutes for electrons radiating at GeV-TeV, in agreement with the minute scale variability observed.
-Therefore, with all the considerations mentioned above, one possible scenario is that the observed flares at VHE rays are produced by freshly accelerated particles via magnetic reconnection with a cooling dominated by SSC cooling.
-We note that the scenario proposed here is based on a magnetically-dominated region.
-However, an alternative picture based on a two-zone model with a weakly magnetised plasma can still offer a good fitting to the SED as long as the blob has a higher Lorentz factor than that of the core, and larger than the expected minimum Lorentz factor obtained from the TeV fast variability, estimated above.
-The most natural way to explain higher Lorentz factor in a small blob compared to the overall blazar zone is magnetic reconnection.
-Nevertheless, and owing to the high degeneracy of blazar SED models (as commented in Sect.
-5), other combinations of parameters can lead to valid representations of the SED and its evolution, possibly relying on non-magnetically-dominated scenarios.
+We find that the SSC cooling timescales of the blob are shorter than the synchrotron timescales, consistent with the Compton dominance of – inferred from the blob SED.
+Table 3 lists the SSC cooling timescales of the blob at .
+For the two nights with intranight variability, our models predict SSC cooling timescales of the order of a few minutes for electrons radiating at GeV-TeV energies, in agreement with the minute-scale variability observed.
+Therefore, considering all of the above, a plausible scenario is that the observed VHE rays flares are produced by freshly accelerated particles via magnetic reconnection with cooling dominated by SSC processes.
+We note that the scenario proposed here assumed a magnetically-dominated region.
+However, an alternative scenario, based on a two-zone model with a weakly magnetised plasma, can still offer a good fitting to the SED, provided that the blob has a higher Lorentz factor than the core, and exceeds the expected minimum Lorentz factor obtained from the TeV fast variability estimated above.
+The most natural explanation for a higher Lorentz factor in the small blob relative to the overall blazar zone is magnetic reconnection.
+Nevertheless, owing to the high degeneracy of blazar SED models (as noted in Sect.
+5), other combinations of parameters can yield valid representations of the SED and its evolution, possibly relying on non-magnetically dominated scenarios.
 7 Conclusions
-BL Lac has shown one of its three brightest -ray and multi-wavelength flares ever between September and November 2023.
-We have characterised its HE and VHE emission, extending up to a few TeV during the brightest periods of the flare.
-One of the key features of this event is the significant remarkably fast VHE -ray flux variability, with doubling flux timescales as short as 8.3 minutes, which sets tight constraints to the size of the VHE emitting region.
-This is also among the fastest flux variations ever detected at VHEs from blazars.
-The HE-VHE spectrum shows a shift of the peak towards higher energies at the highest stages of the flare.
-Accompanying this behaviour we also report a softer-when-brighter evolution of the X-ray spectrum due to a clear transition of the X-ray emission from IC-dominated to synchrotron-dominated, meaning that the source also transitions from LSP to ISP, as seen in other periods for BL Lac in the past (Nilsson et al., 2018).
-The broadband emission was modelled within a leptonic scenario, motivated by recent results reported by IXPE (Middei et al., 2023; Peirson et al., 2023; Agudo et al., 2025; Liodakis et al., 2025).
-One-zone models were not successful when trying to explain the broadband emission.
-Therefore, a co-spatial two-zone SSC model incorporating also the dusty torus EC contribution was used.
-Within this scenario, the evolution of the jet is driven by changes on the electron population of both regions, with injection of freshly accelerated electrons being the cause of the brightest outbursts.
-The minute-scale variability observed constrains the size of the compact blob region to be cm for the adopted Doppler factor , consistent with the value used here.
-This blob region is therefore the responsible of the intranight -ray variability, that can be associated with a plasmoid under a magnetic reconnection event.
+BL Lac exhibited one of its three brightest -ray and multi-wavelength flares on record between September and November 2022.
+We characterised its HE and VHE emission, which extended up to a few TeV during the flares’ brightest periods.
+A key feature of this event is the remarkably fast VHE -ray flux variability, with flux doubling timescales as short as 8.3 min, which sets tight constraints on the size of the VHE emitting region.
+This represents one of the fastest flux variations ever detected at VHEs from blazars.
+The HE-VHE spectrum shows a shift in the peak towards higher energies during the most intense stages of the flare.
+Accompanying this behaviour, we report a softer-when-brighter evolution of the X-ray spectrum resulting from a clear transition of the X-ray emission from IC-dominated to synchrotron-dominated.
+Consequently, the source transitions from LSP to ISP, as previously observed in BL Lac (Nilsson et al., 2018).
+We modelled broadband emission within a leptonic scenario, motivated by recent IXPE results (Middei et al., 2023; Peirson et al., 2023; Agudo et al., 2025; Liodakis et al., 2025).
+One-zone models were unsuccessful in explaining the broadband emission.
+Therefore, we employed a co-spatial two-zone SSC model, also incorporating the EC contribution from the dusty torus.
+Within this scenario, the jet evolution is driven by changes in the electron populations of both regions, with injections of freshly accelerated electrons triggering the brightest outbursts.
+The observed minute-scale variability constrains the size of the compact blob region to cm for the adopted Doppler factor, , consistent with the value used in our model.
+This blob region is therefore responsible for the intranight -ray variability, which can be associated with plasmoids forming during a magnetic reconnection event.
 Appendix A Night-wise LST-1 data analysis results